### v3

# What Matters When Repurposing Diffusion Models for General Dense Perception Tasks?

Abstract
Extensive pre-training with large data is indispensable for downstream geometry and semantic visual perception tasks. Thanks to large-scale text-to-image (T2I) pretraining, recent works show promising results by simply fine-tuning T2I diffusion models for dense perception tasks. However, several crucial design decisions in this process still lack comprehensive justification, encompassing the necessity of the multi-step stochastic diffusion mechanism, training strategy, inference ensemble strategy, and fine-tuning data quality. In this work, we conduct a thorough investigation into critical factors that affect transfer efficiency and performance when using diffusion priors. Our key findings are: 1) High-quality fine-tuning data is paramount for both semantic and geometry perception tasks. 2) The stochastic nature of diffusion models has a slightly negative impact on deterministic visual perception tasks. 3) Apart from fine-tuning the diffusion model with only latent space supervision, task-specific image-level supervision is beneficial to enhance fine-grained details. These observations culminate in the development of GenPercept, an effective deterministic one-step fine-tuning paradigm tailed for dense visual perception tasks. Different from the previous multi-step methods, our paradigm has a much faster inference speed, and can be seamlessly integrated with customized perception decoders and loss functions for image-level supervision, which is critical to improving the fine-grained details of predictions. Comprehensive experiments on diverse dense visual perceptual tasks, including monocular depth estimation, surface normal estimation, image segmentation, and matting, are performed to demonstrate the remarkable adaptability and effectiveness of our proposed method. Code is at:
https://github.com/aim-uofa/GenPercept/
1 Introduction
Recent studies have explored the transferability of text-to-image (T2I) diffusion models to dense visual perception tasks, such as geometry estimation (Ke et al., 2024; Lee et al., 2024; Fu et al., 2024b; Gui et al., 2024; Ye et al., 2024), image segmentation (Van Gansbeke & De Brabandere, 2024; Lee et al., 2024), and inverse rendering (Chen et al., 2024; Kocsis et al., 2024; Zeng et al., 2024). While these works have demonstrated impressive results by repurposing diffusion models for estimating fine-grained geometric and semantic dense prediction maps, the critical design choices made in transferring diffusion models to other dense perception tasks still lack comprehensive justification. This makes it challenging to determine the optimal strategy for achieving the best performance.
For example, Ke et al. (2024) align the visual perception process with the denoising process of Stable Diffusion by fine-tuning all U-Net parameters. They highlight the significance of ”multi-resolution noise” in the forward diffusion process during training, aiming to obtain clean predictions by gradually removing Gaussian noise. On the other hand, (Lee et al., 2024) modify the forward diffusion process by interpolating perception annotations with RGB images instead of using Gaussian noise, and only train the low-rank adaptation (LoRA) Hu et al. (2022) parameters while keeping the U-Net frozen. To the best of our knowledge, the effective components of these approaches have not been thoroughly investigated, leading to uncertainty about which aspects are most crucial for success.
In this work, we examine the design space of repurposing diffusion models for dense visual perception tasks, and explore the key question: What factors are important when adapting diffusion models for general dense perception tasks?
To answer this question, we rethink the importance of both fine-tuning protocols and fine-tuning data. From the perspective of fine-tuning protocols, we categorize recent methods into two main groups: stochastic multi-step generation and deterministic multi-step generation. We explore several critical design spaces, including the diffusion mechanism, key architectural components, training methodologies, and data quality. Our key observations are as follows: 1) The stochastic nature of diffusion models somewhat conflicts with the deterministic requirements of perception tasks, and the amount of noise used significantly impacts perception performance. 2) Strict adherence to traditional diffusion processes is unnecessary. Single-step inference provides similar performance with significantly faster execution. 3) High-quality synthetic fine-tuning data is crucial for geometry perception tasks. From the perspective of fine-tuning data quality, we conduct comprehensive dataset ablation studies on both synthetic datasets and real-world datasets.
Based on the aforementioned observations, we propose GenPercept (see fig. 1), a deterministic fine-tuning paradigm featuring a remarkably simple one-step inference pipeline, an optional customized decoder, and an easily adaptable pixel-specific customized loss. We conduct extensive quantitative and qualitative experiments on a wide range of fundamental visual dense perception tasks, including monocular depth estimation, surface normal estimation, image segmentation, and matting to demonstrate the effectiveness and universality of our method.
In conclusion, our contributions can be summarized as follows: 1) We systematically analyze the design space of current fine-tuning protocols, considering both model architecture and dataset selection, through comprehensive ablation studies. 2) Based on these insights, we propose GenPercept, a simple paradigm that harnesses the power of the pre-trained UNet from diffusion models for generalizable dense visual perception tasks.
Figure 1: Comparisons of three different pipelines. Our GenPercept enables one-step inference and supports pixel-wise losses and customized decoders to replace the cumbersome VAE decoder. We also extend GenPercept to five dense perception tasks including monocular depth estimation, surface normal estimation, dichotomous image segmentation, semantic segmentation, and image matting.
2 preliminary
Standard diffusion models (Rombach et al., 2022; Chen et al., 2023; Song et al., 2020; Ho et al., 2020) define a forward process to inject Gaussian noise $\bm{\epsilon}$ into the input data and a reverse denoising process to estimate noises and generate clean images with a denoiser $\bm{v}_{\theta}$. While diffusion models generate images from noise, the visual perception tasks take RGB images $\mathbf{x}$ as input and estimate the visual perception labels $\mathbf{y}$. Aiming at transferring the prior knowledge of pre-trained diffusion models to computer vision tasks, previous works have proposed to align with the diffusion process and reformulate visual perception tasks as a multi-step denoising process, especially on the monocular depth estimation task. We review the main high-level ideas of previous algorithms in this section and offer detailed formulations in the supplementary material.
Recent works mainly explore two alignment choices. For the stochastic multi-step generation (Ke et al., 2024; Fu et al., 2024b; Gui et al., 2024), the perception tasks are reframed as an RGB image-conditioned denoising process. For training, they add random Gaussian noise $\bm{\epsilon}$ to the ground truth latent $\mathbf{z}^{(\mathbf{y})}$ in the forward diffusion process. For inference, the RGB latent $\mathbf{z}^{(\mathbf{x})}$ is concatenated with the noisy depth latent $\hat{\mathbf{z}}^{(\mathbf{y})}_{t}$ before being sent to U-Net $\bm{v}_{\theta}$ for the $t$-th denoising timestep. Clean depth latent $\hat{\mathbf{z}}^{(\mathbf{y})}_{0}$ will be available after denoising T times. The inference denoising process and forward diffusion process are shown in fig. 1 (a) and fig. 2 (a) separately. These methods perform well but face two significant drawbacks. Firstly, its stochastic nature conflicts with the deterministic nature of perception tasks, which may have a negative impact on performance. Secondly, the high-computation ensemble strategy poses challenges across various perception tasks.
To integrate the deterministic requirements of visual perception tasks with the stochastic diffusion models, deterministic multi-step generation methods such as DMP (Lee et al., 2024) reformulate the forward diffusion process as a blending interpolation between RGB images and ground-truth labels. The inference denoising process and training forward diffusion process are shown in fig. 1 (b) and fig. 2 (a), respectively. Although it addresses the uncertainty inherent in the process, the persistence of RGB texture in the generated perception predictions can detrimentally impact accuracy when visualizing the estimated outcomes. This phenomenon is particularly evident in tasks such as surface normal estimation, as exemplified in fig. 2 (c).
Figure 2: Visualization of different noise forms and proportions in the forward diffusion process. (a) Compared to Gaussian Noise, interpolating RGB images with ground-truth labels can bring less image nature loss. (b) The relationship between the proportion of noise $\sqrt{\bar{\alpha}_{t}}$ and the ($\beta_{start}$, $\beta_{end}$) hyperparameters. (Please see the supplementary for detailed formulations.) (c) Low $\beta$ values during the training of deterministic multi-step generation will lead to the retention of RGB information. Increasing $\beta$ enhances the training difficulty and addresses this issue.
3 Diffusion Models for Visual Perception Tasks
In this section, we explore the necessity and highlight the findings of the multi-step stochastic diffusion mechanism, prior knowledge hides behind the architectural components, training strategy, and fine-tuning data quality. The default experimental setting is similar to Ke et al. (2024) and can be found in supplementary material. We select stochastic method Marigold (Ke et al., 2024) and the deterministic method DMP (Lee et al., 2024) as our baseline.
3.1 The Form and Proportion of Noise in the Forward Diffusion Process
In our initial attempt to integrate the core concepts of Marigold and DMP within a unified codebase, we observed that the performance of the ”baseline DMP” was significantly inferior to that of the ”baseline Marigold.” This discrepancy led us to uncover two notable phenomena: 1) Although deterministic methods are theoretically expected to outperform stochastic, multi-step generation processes for visual perception tasks due to their reduced dependency on random Gaussian noise, our findings contradicted this assumption, with deterministic methods showing inferior performance. 2) The ”multi-resolution noise” strategy introduced by Marigold substantially enhances accuracy.
To further investigate these phenomena, we hypothesize that the challenge associated with denoising from Gaussian noisy latents is substantially more difficult than that from RGB-blended noisy latents. These two different blending paradigms are illustrated in fig. 2(a). We argue that the Gaussian noisy latent less closely approximates natural images and potentially increases the complexity of training, which can enhance the training difficulty. To prove it, we modified the noise blending proportions for both methods by adjusting the $\beta$ values within the DDPM scheduler, effectively increasing the noise component during the forward diffusion process, as depicted in fig. 2(b).
Our quantitative and qualitative analyses, presented in table 1 and fig. 2(c), indicate that increasing training challenges lead to improved model performance. As the $\beta$ value increases, the impact of ”multi-resolution noise” diminishes. However, unlike DMP, which continues to show performance improvements, Marigold’s performance begins to be slightly unstable but shows a rough improvement trend when the $\beta$ value is sufficiently high. We consider it brought by the randomness of the Gaussian Noise, which dominates most components of the input latent.
When $\beta$ reaches 1, the noisy label latent input effectively becomes a pure noise latent, akin to the RGB latent in DMP. Consequently, the target output of the ”v-prediction” diffusion model shifts to the negative value of the ground truth latent, eliminating any intermediate state of blending between RGB noise latent and ground-truth latent. The denoising process becomes repeats of “predict target label $\hat{\mathbf{z}}^{(\mathbf{y})}_{0}$ from RGB image $\mathbf{z}^{(\mathbf{x})}$ and add (RGB) noise back to the RGB image $\mathbf{z}^{(\mathbf{x})}$”, whose estimation results in any timestep $t$ are exactly identical. Therefore, we propose to reduce the DDIM steps to one and call it ”deterministic one-step perception”. The resulting inference speeds are significantly faster, with performance remaining largely consistent. We consider it as “our baseline” for the subsequent analyses.
Table 1: Comprehensive quantitative comparisons about the impact of noise forms and proportions in the forward diffusion process on monocular depth estimation. Visualizations of different noise forms and the effect of $\beta$ values are shown in fig. 2. The performance of DMP improves steadily, while Marigold shows initial improvements followed by a decline. When $\beta$ reaches 1, the inference process can be reduced to one step without compromising performance. “Rank” means the average rank of ten evaluation performance (smaller is better).
Type Noise Multi-res Steps ($\beta_{start}$, $\beta_{end}$) KITTI NYU ScanNet DIODE ETH3D Rank
[…]
Marigold Gaussian $\checkmark$ 10 (0.0002125, 0.003) 0.358 0.462 0.297 0.555 0.246 0.625 0.494 0.565 0.267 0.640 7.0
Marigold Gaussian $\checkmark$ 10 (0.000425, 0.006) 0.122 0.854 0.106 0.887 0.136 0.829 0.345 0.716 0.086 0.927 5.8
baseline Marigold Gaussian $\checkmark$ 10 (0.00085, 0.012) 0.099 0.909 0.063 0.956 0.075 0.937 0.316 0.764 0.075 0.947 3.6
Marigold Gaussian $\checkmark$ 10 (0.0034, 0.048) 0.100 0.906 0.057 0.963 0.063 0.957 0.308 0.768 0.074 0.948 2.3
Marigold Gaussian $\checkmark$ 10 (0.1360, 0.192) 0.119 0.861 0.058 0.963 0.061 0.961 0.315 0.760 0.073 0.950 2.8
Marigold Gaussian $\checkmark$ 10 (0.5440, 0.768) 0.124 0.852 0.060 0.961 0.064 0.958 0.322 0.749 0.079 0.943 4.7
Marigold Gaussian $\checkmark$ 10 (1.0, 1.0) 0.104 0.897 0.055 0.965 0.059 0.962 0.312 0.762 0.069 0.955 1.7
Marigold Gaussian $\checkmark$ 1 (1.0, 1.0) 0.104 0.897 0.055 0.965 0.059 0.962 0.312 0.762 0.069 0.955 -
Marigold Gaussian $\times$ 10 (0.0002125, 0.003) 0.587 0.255 0.337 0.490 0.257 0.604 0.600 0.469 0.372 0.503 7.0
Marigold Gaussian $\times$ 10 (0.000425, 0.006) 0.536 0.289 0.313 0.527 0.248 0.621 0.565 0.499 0.328 0.575 6.0
baseline Marigold Gaussian $\times$ 10 (0.00085, 0.012) 0.153 0.807 0.162 0.802 0.187 0.737 0.411 0.641 0.157 0.826 5.0
Marigold Gaussian $\times$ 10 (0.0034, 0.048) 0.101 0.907 0.058 0.963 0.066 0.954 0.309 0.765 0.074 0.950 2.4
Marigold Gaussian $\times$ 10 (0.1360, 0.192) 0.115 0.870 0.056 0.965 0.060 0.961 0.313 0.763 0.072 0.953 2.3
Marigold Gaussian $\times$ 10 (0.5440, 0.768) 0.124 0.848 0.059 0.963 0.063 0.958 0.318 0.752 0.077 0.946 3.7
Marigold Gaussian $\times$ 10 (1.0, 1.0) 0.102 0.901 0.054 0.966 0.059 0.962 0.312 0.762 0.071 0.955 1.5
Marigold Gaussian $\times$ 1 (1.0, 1.0) 0.102 0.901 0.054 0.966 0.059 0.962 0.312 0.762 0.071 0.955 -
DMP RGB $\times$ 10 (0.0002125, 0.003) 0.476 0.336 0.267 0.601 0.216 0.677 0.457 0.588 0.185 0.757 6.9
DMP RGB $\times$ 10 (0.000425, 0.006) 0.265 0.630 0.201 0.072 0.195 0.717 0.386 0.674 0.116 0.880 6.1
baseline DMP RGB $\times$ 10 (0.00085, 0.012) 0.134 0.837 0.117 0.871 0.147 0.808 0.353 0.721 0.093 0.919 5.0
DMP RGB $\times$ 10 (0.0034, 0.048) 0.107 0.890 0.077 0.939 0.087 0.923 0.318 0.766 0.078 0.940 3.8
DMP RGB $\times$ 10 (0.1360, 0.192) 0.107 0.890 0.063 0.959 0.068 0.955 0.305 0.773 0.073 0.948 2.2
DMP RGB $\times$ 10 (0.5440, 0.768) 0.106 0.897 0.061 0.959 0.066 0.952 0.309 0.768 0.075 0.945 2.3
DMP RGB $\times$ 10 (1.0, 1.0) 0.100 0.902 0.053 0.966 0.059 0.961 0.309 0.768 0.068 0.956 1.2
Our baseline RGB $\times$ 1 (1.0, 1.0) 0.100 0.902 0.053 0.966 0.059 0.961 0.309 0.768 0.068 0.956 -
Finding 1.
The inherent stochasticity of diffusion models can lead to minor, unstable performance fluctuations in deterministic perception tasks. By increasing the noise proportion, the multi-step generation can be simplified to a one-step fine-tuning paradigm without any loss of performance in both stochastic and deterministic methods.
3.2 Where Does the Rich Visual Knowledge Reside in Diffusion Models?
Based on the baseline we proposed in §3.1, we conduct detailed ablation studies to thoroughly investigate the necessity of each component of Stable Diffusion. Results are reported in table 2.
Table 2: Explorations on the impact of the Stable Diffusion components on depth estimation. Customized decoders and losses can also enable inference acceleration and performance improvement.
Setting Loss KITTI NYU ScanNet DIODE ETH3D
[…]
Our baseline MSE (Latent) 0.100 0.902 0.053 0.966 0.059 0.961 0.309 0.768 0.068 0.956
Train U-Net from scratch MSE (Latent) 0.219 0.650 0.186 0.736 0.183 0.729 0.426 0.614 0.185 0.741
Train VAE decoder from scratch MSE (Image) 0.096 0.916 0.055 0.964 0.058 0.964 0.302 0.759 0.071 0.950
Baseline + Image MSE loss MSE (Image) 0.097 0.915 0.054 0.964 0.059 0.964 0.305 0.760 0.071 0.953
Baseline + Image customized loss MSE + SSI + 0.094 0.923 0.052 0.966 0.056 0.965 0.302 0.767 0.066 0.967
Grad. (Image)
Train DPT decoder from scratch MSE (Image) 0.099 0.912 0.055 0.964 0.058 0.963 0.302 0.759 0.069 0.956
Denoiser. We reinitialize the U-Net parameter and train it from scratch on the same datasets. Without prior knowledge of large data from LAION-5B, the network performs poorly and loses the generalization. This indicates that most of the prior knowledge is stored in the denoiser module.
VAE AutoEncoder. The VAE encoder’s original architecture is kept intact to maintain the consistency of the encoding process. For the VAE decoder, we train it from scratch with image pixel MSE loss. Without pre-trained parameters of the VAE decoder, it still performs well.
Customized Head and Loss. The deterministic one-step perception pipeline enables customized heads and loss functions for training. By utilizing a DPT decoder or the loss functions of DepthAnythingv2 (Yang et al., 2024b), we can implement a lightweight decoder that supervises pixel-wise information at a higher resolution rather than latent features at 1/8 resolution. This approach can not only accelerate inference times but also enhance the acquisition of fine-grained details.
Finding 2.
The primary perceptual prior knowledge of diffusion models is encapsulated within the denoiser. Customized heads and loss functions can achieve enhanced inference speed and fine-grained details.
3.3 What About the Timesteps and Text Prompts?
The timesteps and text prompts are crucial elements in utilizing the Stable Diffusion model to generate diverse images. We conducted ablation studies to investigate their significance. The results reported in table 3 indicate a negligible difference between various settings. Owing to the inherent certainty associated with visual perception tasks, the diversity typically offered by the textual inputs appears to be unnecessary. Similarly, the utility of timesteps is reduced, as the single-step paradigm does not require progressive denoising.
Table 3: Quantitative comparisons among different timesteps and text prompts of diffusion models on depth estimation.
Setting Text Prompt Train / Infer KITTI NYU ScanNet DIODE ETH3D
[…]
Our baseline “” Random / 1 0.100 0.902 0.053 0.966 0.059 0.961 0.309 0.768 0.068 0.956
Valid text input “A high quality RGB image” Random / 1 0.101 0.900 0.053 0.967 0.058 0.964 0.312 0.762 0.070 0.954
Random text input “F3@qV!k2*#Zp^n%1Lz” Random / 1 0.099 0.904 0.054 0.965 0.059 0.963 0.311 0.763 0.069 0.955
Timestep1 “” 1 / 1 0.100 0.906 0.054 0.965 0.060 0.961 0.304 0.769 0.069 0.956
Timestep500 “” 500 / 500 0.102 0.897 0.053 0.966 0.059 0.961 0.307 0.765 0.068 0.956
Timestep900 “” 900 / 900 0.105 0.891 0.054 0.966 0.058 0.964 0.309 0.762 0.068 0.953
Finding 3.
The timesteps and text prompts of diffusion models are negligible for the performance of visual perception tasks.
3.4 How to Leverage the U-Net’s Prior Knowledge?
The significance of the denoiser cannot be overstated. However, the strategies for its utilization are worth careful consideration. Should we freeze the denoiser, utilize its intermediate features, and merely fine-tune the decoder for specific tasks? Alternatively, could we employ LoRA (Hu et al., 2022) instead of extensively fine-tuning the entire denoiser? Unfortunately, the evidence suggests that neither approach is ideal. As illustrated in table 4, freezing the denoiser significantly compromises performance. Although incorporating LoRA offers some advantages, it can not fully leverage the potential of denoiser, especially with regular LoRA ranks of 4 and 16. This limitation likely stems from the substantial differences between the noise-to-image denoising process and the image-to-perception prediction task.
Table 4: Explorations on the paradigms to leverage U-Net’s prior knowledge on depth estimation.
Setting LoRA KITTI NYU ScanNet DIODE ETH3D
[…]
Our baseline - 0.100 0.902 0.053 0.966 0.059 0.961 0.309 0.768 0.068 0.956
Freeze U-Net + Train DPT decoder - 0.144 0.803 0.086 0.931 0.097 0.911 0.309 0.768 0.068 0.956
Train U-Net with LoRA 4 0.211 0.644 0.095 0.914 0.100 0.902 0.372 0.689 0.121 0.864
Train U-Net with LoRA 16 0.166 0.746 0.085 0.931 0.087 0.927 0.352 0.712 0.104 0.901
Train U-Net with LoRA 64 0.138 0.817 0.077 0.944 0.079 0.940 0.336 0.734 0.089 0.930
Train U-Net with LoRA 256 0.133 0.827 0.069 0.952 0.073 0.947 0.325 0.745 0.088 0.933
Train U-Net with LoRA 1024 0.125 0.849 0.067 0.955 0.074 0.947 0.324 0.747 0.084 0.939
Finding 4.
Fine-tuning the denoiser is preferable compared to either merely utilizing its intermediate features or training it with LoRA.
3.5 Is the Training Data Quality Essential?
The quality of annotations in real datasets is often lower compared to synthetic datasets, where data is precisely rendered via simulators. In table 5, we explore the impact of data quality on the fine-tuning process. We sample the same distribution of real data, consisting of 90% from approximately 50K indoor images from the Taskonomy dataset (Zamir et al., 2018) and 10% from about 40K outdoor images from the Cityscapes dataset (Cordts et al., 2016). With lower annotation quality, the model achieves slightly worse performance. Also, the visualization in the supplementary material indicates that noisy data significantly influences detailed predictions in visual perception tasks.
Table 5: Investigations into the impact of training data quality on depth estimation.
Data Quality Datasets KITTI NYU ScanNet DIODE ETH3D
[…]
Synthetic Data Hypersim (50K) + Virtual KITTI (40K) 0.100 0.902 0.053 0.966 0.059 0.961 0.309 0.768 0.068 0.956
Real Data Taskonomy (50K) + Cityscapes (40K) 0.123 0.857 0.055 0.966 0.062 0.958 0.293 0.762 0.074 0.947
Finding 5.
Data quality affects the fine-grained details of dense predictions significantly.
3.6 Summary of the obervations
Based on the preceding analysis, an effective approach to leveraging the prior knowledge of diffusion models is to use them as single-step deterministic perception estimators. This can be done with either a VAE decoder or a customized lightweight decoder. Additionally, employing pixel-specific customized loss functions can further enhance detail and overall performance. We compare our deterministic single-step perception method with previous multi-step paradigms in fig. 1. In the following section, we extend these findings to a broader set of visual perception tasks, including surface normal estimation, semantic image segmentation, dichotomous image segmentation, and image matting.
4 Experiments on Various Dense Visual Perceptual Tasks
In this section, we empirically show the robust transfer ability of our GenPercept on diverse visual tasks. Unless specified otherwise, we freeze the VAE AutoEncoder and fine-tune the U-Net of Stable Diffusion v2.1 to estimate the ground-truth label latent for 30000 iterations, with a resolution of (768, 768), a batch size of 32, and a learning rate of 3e-5. Different customized loss functions are utilized to improve the performance further on dense visual perception tasks.
4.1 Geometric Estimation
[…]
Table 6: Quantitative comparison of affine-invariant depth estimation on five zero-shot datasets.
Method Training KITTI NYU ScanNet DIODE ETH3D
[…]
MiDaS (Ranftl et al., 2020) 2M 0.236 0.630 0.111 0.885 0.121 0.846 0.332 0.715 0.184 0.752
Omnidata (Eftekhar et al., 2021) 12.2M 0.149 0.835 0.074 0.945 0.075 0.936 0.339 0.742 0.166 0.778
DPT-large (Ranftl et al., 2021) 1.4M 0.100 0.901 0.098 0.903 0.082 0.934 0.182 0.758 0.078 0.946
DepthAnything (Yang et al., 2024a) 63.5M 0.080 0.946 0.043 0.980 0.043 0.981 0.261 0.759 0.058 0.984
DepthAnything v2 (Yang et al., 2024b) 62.6M 0.080 0.943 0.043 0.979 0.042 0.979 0.321 0.758 0.066 0.983
Metric3D v2 (Hu et al., 2024) 16M 0.052 0.979 0.039 0.979 0.023 0.989 0.147 0.892 0.040 0.983
DiverseDepth (Yin et al., 2020) 320K 0.190 0.704 0.117 0.875 0.109 0.882 0.376 0.631 0.228 0.694
LeReS (Yin et al., 2021) 354K 0.149 0.784 0.090 0.916 0.091 0.917 0.271 0.766 0.171 0.777
HDN (Zhang et al., 2022a) 300K 0.115 0.867 0.069 0.948 0.080 0.939 0.246 0.780 0.121 0.833
Marigold (Ke et al., 2024) 74K 0.099 0.916 0.055 0.964 0.064 0.951 0.308 0.773 0.065 0.960
GeoWizard (Fu et al., 2024b) 280K 0.129 0.851 0.059 0.959 0.066 0.953 0.328 0.753 0.077 0.940
DepthFM (Gui et al., 2024) 63K 0.174 0.718 0.082 0.932 0.095 0.903 0.334 0.729 0.101 0.902
Our GenPercept (Depth) 90K 0.094 0.923 0.052 0.966 0.056 0.965 0.302 0.767 0.066 0.957
Our GenPercept (Disparity) 90K 0.080 0.934 0.058 0.969 0.063 0.960 0.226 0.741 0.096 0.959
Our GenPercept (Disparity + DPT head) 90K 0.078 0.935 0.059 0.967 0.064 0.961 0.228 0.740 0.094 0.961
Table 7: Quantitative comparison of surface normal estimation on three zero-shot datasets. We evaluate mean $\downarrow$, median error $\downarrow$ (med.), and the percentages of pixels $\uparrow$ with five thresholds.
Method Training NYU v2 ScanNet Sintel
Samples mean med. $5.0^{\circ}$ $7.5^{\circ}$ $11.25^{\circ}$ $22.5^{\circ}$ $30^{\circ}$ mean med. $5.0^{\circ}$ $7.5^{\circ}$ $11.25^{\circ}$ $22.5^{\circ}$ $30^{\circ}$ mean med. $5.0^{\circ}$ $7.5^{\circ}$ $11.25^{\circ}$ $22.5^{\circ}$ $30^{\circ}$
Omnidata v1 (Eftekhar et al., 2021) 12.2M 23.1 12.9 21.6 33.4 45.8 66.3 73.6 22.9 12.3 21.5 34.5 47.4 66.1 73.2 41.5 35.7 3.0 5.8 11.4 30.4 42.0
Ominidata v2 (Kar et al., 2022) 12.2M 17.2 9.7 25.3 40.2 55.5 76.5 83.0 16.2 8.5 29.1 44.9 60.2 79.5 84.7 40.5 35.1 4.6 7.9 14.7 33.0 43.5
Metric3D v2 (Hu et al., 2024) 8.8M 13.5 6.7 40.1 53.5 65.9 82.6 87.7 11.8 5.5 46.6 60.7 71.6 85.4 89.7 22.8 14.2 18.4 28.5 41.6 66.7 75.8
DINSE (Bae & Davison, 2024a) 160K 16.4 8.4 32.8 46.3 59.6 77.7 83.5 16.2 8.3 29.8 45.9 61.0 78.7 84.4 34.9 28.1 8.9 14.1 21.5 41.5 52.7
Geowizard (Fu et al., 2024b) 280K 19.8 11.2 18.0 32.7 50.2 73.0 79.9 21.1 11.9 15.9 29.7 47.4 70.7 77.8 36.1 28.4 4.1 8.6 16.9 39.8 52.5
Our GenPercept (Latent MSE loss) 90K 17.4 9.5 23.3 40.0 56.3 76.8 83.0 16.3 8.9 25.8 42.7 59.6 79.4 84.8 44.4 31.6 3.4 7.5 15.0 37.0 48.0
Our GenPercept (Image angular loss) 90K 16.4 8.0 33.3 47.8 60.9 78.3 83.7 15.2 7.4 33.9 50.7 65.0 80.9 85.7 34.6 26.2 5.2 9.8 18.4 43.8 55.8
Surface Normal Estimation. The surface normal estimation aims to predict a vector perpendicular to tangent plane of the surface at each point P, which represents the orientation of the object’s surface. For evaluation, we compute the angular error on three zero-shot datasets including NYU (Silberman et al., 2012), ScanNet (Dai et al., 2017), and Sintel (Butler et al., 2012). The mean $\downarrow$, median $\downarrow$, and the percentages of pixels $\uparrow$ with error below thresholds [5, 7.5, 11.25, 22.5, 30] are reported. Invalid regions are filtered out and the metrics are averaged on all images.
Quantitative Evaluation. Quantitative results on monocular depth estimation and surface normal estimation are shown in table 6 and table 7, respectively. Even trained on limited synthetic datasets only, our GenPercept shows much robustness and achieves promising performance on diverse unseen scenes. For monocular depth models, we train them with pixel-wise MSE loss, scale-shift-invariant loss (Ranftl et al., 2020), and gradient loss (Ranftl et al., 2020). Furthermore, our disparity model (inverse of the depth) shows much better performance on datasets with outdoor scenes, such as KITTI and DIODE, but less performance on indoor datasets. By replacing the VAE decoder with a lightweight DPT head, GenPercept can infer faster without bearing the performance penalty. For surface normal estimation, the image angular loss brings significant performance improvement thanks to our one-step estimation paradigm.
Qualitative Results. Qualitative visualizations are shown in fig. 3. We observe excellent generalization of our models in that they can estimate accurate geometric information and promising details not only on diverse real scenes and synthetic scenes, but also on comics, color drafts, and even some sketches.
Figure 3: Qualitative results of monocular depth and surface normal estimation. The model works surprisingly well on out-of-domain images (sketch and cartoon images).
4.2 Image Segmentation
Table 8: Quantitative results of dichotomous image segmentation on DIS5K validation and testing sets. Additional cross-dataset evaluation is provided in the supplementary material.
Dataset DIS-VD DIS-TE4 Overall DIS-TE (1-4)
Metric $maxF_{\beta}\uparrow$ $F^{w}_{\beta}\uparrow$ ${}M\downarrow$ $S_{\alpha}\uparrow$ $E_{\phi}^{m}\uparrow$ ${\rm HCE}_{\gamma}\downarrow$ $maxF_{\beta}\uparrow$ $F^{w}_{\beta}\uparrow$ ${}M\downarrow$ $S_{\alpha}\uparrow$ $E_{\phi}^{m}\uparrow$ ${\rm HCE}_{\gamma}\downarrow$ $maxF_{\beta}\uparrow$ $F^{w}_{\beta}\uparrow$ ${}M\downarrow$ $S_{\alpha}\uparrow$ $E_{\phi}^{m}\uparrow$ ${\rm HCE}_{\gamma}\downarrow$
U2Net (Qin et al., 2020) 0.748 0.656 0.090 0.781 0.823 1413 0.795 0.705 0.087 0.807 0.847 3653 0.761 0.670 0.083 0.791 0.835 1333
SINetV2 (Fan et al., 2021a) 0.665 0.584 0.110 0.727 0.798 1568 0.699 0.616 0.113 0.744 0.824 3683 0.693 0.608 0.101 0.747 0.822 1411
HySM (Nirkin et al., 2021) 0.734 0.640 0.096 0.773 0.814 1324 0.782 0.693 0.091 0.802 0.842 3331 0.757 0.665 0.084 0.792 0.834 1218
IS-Net (Qin et al., 2022) 0.791 0.717 0.074 0.813 0.856 1116 0.827 0.753 0.072 0.830 0.870 2888 0.799 0.726 0.070 0.819 0.858 1016
Our Genpercept 0.857 0.835 0.04 0.87 0.934 1511 0.848 0.829 0.049 0.854 0.938 3799 0.863 0.839 0.039 0.872 0.936 1364
Table 9: Quantitative results of semantic segmentation performance on Hypersim and ADE20k datasets.
[…]
(Hypersim) (ADE20K)
Our GenPercept 46.0 34.1
Our GenPercept 52.9 38.3
Table 10: Quantitative comparisons of image matting on the P3M-500-NP test set.
Method SAD $\downarrow$ MAD $\downarrow$ MSE $\downarrow$ CONN $\downarrow$
HATT (Qiao et al., 2020) 30.35 0.0176 0.0072 27.42
SHM (Chen et al., 2018b) 20.77 0.0122 0.0093 17.09
MODNet (Ke et al., 2022) 16.70 0.0097 0.0051 13.81
P3M-Net (Li et al., 2021) 11.23 0.0065 0.0035 12.51
Our GenPercept 12.77 0.0074 0.0027 10.46
Dichotomous Image Segmentation. This is a category-agnostic, high-quality object segmentation task that accurately separates the object from the background in an image. Consistent with previous methods, we use the six evaluation metrics specified in the DIS task, which include maximal F-measure ($maxF_{\beta}\uparrow$(Achanta et al., 2009), weighted F-measure ($F^{w}_{\beta}\uparrow$(Margolin et al., 2014), mean absolute error (${}M\downarrow$(Perazzi et al., 2012), structural measure ($S_{\alpha}\uparrow$(Fan et al., 2017), mean enhanced alignment measure ($E_{\phi}^{m}\uparrow$(Fan et al., 2018; 2021b) and human correction efforts (${\rm HCE}_{\gamma}\downarrow$(Qin et al., 2022). We choose DIS5K (Qin et al., 2022) as the training and testing dataset. We utilize DIS-TR for training and evaluate our model on DIS-VD and DIS-TE subsets. The pixel-wise MSE loss is utilized during training.
Quantitative results of dichotomous image segmentation are shown in table 8, respectively. We only show partial results due to paper page limitations, full comparisons are accessible in the supplementary material. Our proposed model outperforms all other methods by a significant margin on this challenging dataset across most evaluation metrics. As shown in Fig. 5, the segmentation results of our approach are more refined, providing a more detailed foreground mask. For thin lines and meticulous objects that are difficult for previous methods to process, our method can also output accurate segmentation results.
Semantic Image Segmentation. This is a fundamental computer vision task that involves classifying each pixel in an image into a specific category or class. For training, we utilized the indoor synthetic dataset, HyperSim (Roberts et al., 2021), which comprises 40 semantic segmentation class labels. We encode different classes into 3-channel colormaps, treat the task as a regression problem, and fine-tune the original Stable Diffusion with the pixel-wise MSE loss. As demonstrated in fig. 4, the model generalizes well to classes within the HyperSim annotations, such as chairs and desks, but struggles with unrecognized categories such as cats and cars.
Another choice involves using a customized segmentation head. We incorporate a custom segmentation head, namely UperNet (Xiao et al., 2018b), onto the multi-level features extracted by UNet, and the quantitative results are presented in table 10. In addition to evaluating on Hypersim (Roberts et al., 2021), we also test the model’s zero-shot performance on a subset of the ADE20k (Zhou et al., 2017) validation set, which contains overlapping classes. The results demonstrate the model’s capacity for zero-shot generalization.
Figure 4: Quanlitative results of semantic segmentation in the wild. Trained on the synthetic indoor Hypersim dataset, GenPercept shows much robustness on the trained categories of complex in-the-wild images, e.g., yellow chairs, green floor, and light blue wall. Due to the limited annotation categories and little negative label of “unknown category”, it sometimes fails in outdoor scenes and unseen categories such as cats and cars.
Figure 5: Qualitative comparison of dichotomous image segmentation.
4.3 Image Matting
Task Definition. Image matting aims to extract the foreground, background, and alpha mask from an image. Traditional approaches depend on supplementary inputs that delineate foreground, background, and ambiguous areas to reduce uncertainty. Automatic image matting seeks to remove this dependency by directly estimating these components from the image alone.
Implementation Details. We utilize P3M10K (Li et al., 2021), the largest portrait matting dataset with high-resolution portrait images along with high-quality alpha masks to train our model. The training set contains 9,421 high-quality images and annotations and the test set P3M-500-NP is composed of 500 public images from the Internet. We train our GenPercept with pixel-wise MSE loss to further improve the final performance.
Quantitative and Qualitative Results. We evaluate metrics including the sum of absolute differences (SAD), mean squared error (MSE), mean absolute difference (MAD), gradient (Grad.), and Connectivity (Conn.) on the P3M-500-NP test set. SAD and MAD measure the mean L1 distance between predictions and ground truth labels. MSE and CONN focus on L2 distance and connectivity that better reflects human intuition. As shown in table 10, our GenPercept is comparable with the state-of-the-art methods. Quantitative results of image matting are shown in fig. 6.
Figure 6: Visualization of image matting on the P3M-500-NP test set.
5 Related work
Vision Pre-Training. Models pretrained on large-scale datasets possess powerful feature extraction capabilities, enabling them to be effectively transferred to a wide range of visual tasks. For instance, the ResNet (He et al., 2016) model pretrained on ImageNet (Russakovsky et al., 2015) can be fine-tuned and applied to perception tasks. By means of contrastive learning, MoCo (He et al., 2020) and CLIP (Radford et al., 2021) acquire rich visual and semantic representations, leveraging their advantages in joint visual and semantic modeling to enhance the performance of multimodal tasks. DINO (Zhang et al., 2022b), through self-distillation, endows Vision Transformer and convolutional networks with comparable visual representation quality and demonstrates that self-supervised ViT representations contain explicit semantic segmentation information. In our work, we leverage Stable Diffusion (Rombach et al., 2022) as a prior for scene understanding and transfer it to various perception tasks.
Diffusion Priors for Dense Prediction. Several works explore to use the priors of generative models for perceptual tasks. Some works (Bhattad et al., 2024; Du et al., 2023) demonstrate that generative models encode property maps of the scene. By finding latent variable offsets, using LoRA (Hu et al., 2022), etc., generative models can directly produce intrinsic images like surface normals, depth, albedo, etc. LDMSeg (Van Gansbeke & De Brabandere, 2024) devises an image-conditioned sampling process, enabling diffusion models to directly output panoptic segmentation. UniGS (Qi et al., 2023) proposes location-aware color encoding and decoding strategies, allowing diffusion models to support referring segmentation and entity segmentation. Marigold (Ke et al., 2024) fine-tunes diffusion model on limited synthetic data, enabling it to support affine-invariant monocular depth estimation and exhibit strong generalization performance. However, Marigold is time-consuming due to the need for multiple iterations of denoising. Additionally, the Gaussian noise leads to inconsistent results across inferences, requiring aggregation over multiple inferences. Xiang et al. (2023) train a denoising auto-encoder for image classification. The difference of their method compared with traditional denoising auto-encoder is that input images are encoded into a latent code and denoising is performed in the latent space rather than the pixel space. They show good results on very small-scale datasets (CIFAR and ImageNet-tiny) to prove the concept and no results were reported on larger datasets. Furthermore, GeoWizard (Fu et al., 2024a) extends the generative capabilities of Marigold, achieving better performance in joint depth and normal estimation, which enhances applications like 3D reconstruction and novel view synthesis. Moreover, DepthFM (Gui et al., 2024) addresses the speed challenge of Marigold by employing flow matching, offering a fast and efficient monocular depth estimation model.
6 Conclusion
In this work, we introduce GenPercept, an embarrassingly straightforward yet powerful approach to re-use the off-the-shelf UNet trained using diffusion processes. In essence, we initialize our image perception networks using pre-trained diffusion models trained on the extensive LAION dataset, without making many modifications. GenPercept demonstrates the capability to effectively leverage pre-trained diffusion models across a range of downstream dense perception tasks. We contend that our proposed methodology provides an efficient and potent paradigm for harnessing the capabilities of pre-trained diffusion models in dense visual perception tasks.
Supplementary Materials of
What Matters When Repurposing Diffusion Models for General Dense Perception Tasks?
Appendix A The formulation of Three Different Pipelines
A.1 Stochastic Multi-Step Generation
For the training process, the RGB image $\mathbf{x}$ and ground-truth label $\mathbf{y}$ are encoded into the latent space with the VAE encoder $\mathbf{z}^{(\mathbf{x})}=\mathcal{E}(\mathbf{x})$, $\mathbf{z}^{(\mathbf{y})}=\mathcal{E}(\mathbf{y})$. The Gaussian noise $\bm{\epsilon}$ is added to the ground-truth label latent $\mathbf{z}^{(\mathbf{y})}$, and the noisy label latent $\mathbf{z}^{(\mathbf{y})}_{t}$ is concatenated with the clean image latent $\mathbf{z}^{(\mathbf{x})}$ as U-Net input $\mathbf{z}_{t}$ for each timestep:
$\begin{gathered}\mathbf{z}^{(\mathbf{y})}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{z}^{(\mathbf{y})}+\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon},~{}~{}t=[1,...,T],\\ \mathbf{z}_{t}=\text{concat}(\mathbf{z}^{(\mathbf{y})}_{t},\mathbf{z}^{(\mathbf{x})}),\end{gathered}$ (1)
where $\bar{\alpha}_{t}=\prod_{s=1}^{t}{(1-\beta_{s})}$, and $\beta_{s}$ is sampled from a variance schedule $\{\beta_{t}\in(0,1)\}_{t=1}^{T}$. Then, the denoiser $\bm{v}_{\theta}(\cdot,\cdot)$ is enforced to learn the velocity of “v-prediction” Salimans & Ho (2021) from a timestep $t$ and the corresponding input $\mathbf{z}_{t}$. During training, the parameters of VAE are frozen, and only the denoiser $\bm{v}_{\theta}$ is fine-tuned. The timestep is uniformly sampled from 1 to $T$.
$\begin{gathered}\mathcal{L}=\mathbb{E}_{\mathbf{z}^{(\mathbf{y})},\bm{\epsilon}\sim\mathcal{N}(0,I),t\sim\mathcal{U}(T)}\left\|(\sqrt{\bar{\alpha}_{t}}\bm{\epsilon}-\sqrt{1-\bar{\alpha}_{t}}\mathbf{z}^{(\mathbf{y})})-\bm{v}_{\theta}(\mathbf{z}_{t},t)\right\|^{2}_{2}.\end{gathered}$ (2)
For inference, a Gaussian noise $\bm{\epsilon}_{t}$ is randomly sampled and denoised step by step with the denoiser $\bm{v}_{\theta}(\cdot,\cdot)$.
$\begin{gathered}\hat{\mathbf{z}}^{(\mathbf{y})}_{T}=\bm{\epsilon},~{}~{}\mathbf{z}_{t}=\text{concat}(\hat{\mathbf{z}}^{(\mathbf{y})}_{t},\mathbf{z}^{(\mathbf{x})}),~{}~{}t=[T,...,1],\\ \hat{\mathbf{z}}^{(\mathbf{y})}_{t\rightarrow 0}=\sqrt{\bar{\alpha}_{t}}\cdot\hat{\mathbf{z}}^{(\mathbf{y})}_{t}-\sqrt{1-\bar{\alpha}_{t}}\cdot\bm{v}_{\theta}(\mathbf{z}_{t},t),~{}~{}\hat{\bm{\bm{\epsilon}}}_{t}=\sqrt{\bar{\alpha}_{t}}\cdot\bm{v}_{\theta}(\mathbf{z}_{t},t)-\sqrt{1-\bar{\alpha}_{t}}\cdot\hat{\mathbf{z}}^{(\mathbf{y})}_{t},\\ \hat{\mathbf{z}}^{(\mathbf{y})}_{t-1}=\sqrt{\bar{\alpha}_{t-1}}\cdot\hat{\mathbf{z}}^{(\mathbf{y})}_{t\rightarrow 0}+\sqrt{1-\bar{\alpha}_{t-1}}\cdot\hat{\bm{\bm{\epsilon}}}_{t},~{}~{}\hat{\mathbf{y}}=\mathcal{D}(\hat{\mathbf{z}}^{(\mathbf{y})}_{1\rightarrow 0}).\end{gathered}$ (3)
where the denoising process first computes the estimated noise $\hat{\bm{\bm{\epsilon}}}_{t}$ and the predicted clean latent code $\hat{\mathbf{z}}^{(\mathbf{y})}_{t\rightarrow 0}$ of timestep $t$ from the current latent code $\hat{\mathbf{z}}^{(\mathbf{y})}_{t}$ and the predicted velocity $\bm{v}_{\theta}(\mathbf{z}_{t},t)$. Then, it adds the computed noise $\hat{\bm{\bm{\epsilon}}}_{t}$ back to $\hat{\mathbf{z}}^{(\mathbf{y})}_{t\rightarrow 0}$ to get the latent code of timestep $t-1$. After repeating it for $T$ times, the predicted clean latent code $\hat{\mathbf{z}}^{(\mathbf{y})}_{1\rightarrow 0}$ is computed and sent to the VAE decoder $\mathcal{D}$ and estimate the target label $\hat{\mathbf{y}}$. During inference, $m$ randomly sampled noises are introduced to estimate $m$ different predictions, and they are averaged with an ensemble process to reduce the randomness of prediction and improve performance.
A.2 Deterministic Multi-Step Generation
The inherent random nature of diffusion models makes them challenging to apply to perceptual tasks, which typically aim for accurate results. As a result, existing works have replaced the original Gaussian noise with the target image as RGB noise (Bansal et al., 2024; Lee et al., 2024). Technically, rather than introducing the random Gaussian noise $\bm{\epsilon}$ , we blend the ground-truth label latent $\mathbf{z}^{(\mathbf{y})}=\mathcal{E}(\mathbf{y})$ with the RGB image latent $\mathbf{z}^{(\mathbf{x})}=\mathcal{E}(\mathbf{x})$, which is formulated as:
$\begin{gathered}\mathbf{z}_{t}=\mathbf{z}^{(\mathbf{y})}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{z}^{(\mathbf{y})}+\sqrt{1-\bar{\alpha}_{t}}\mathbf{z}^{(\mathbf{x})},~{}~{}t=[1,...,T],\\ \end{gathered}$ (4)
Figure 7: Qualitative visualization of ablation study.
Figure 8: Visualization of different noise forms and proportions in the forward diffusion process.
Furthermore, the input latent has been modified to the latent code of input image $\mathbf{z}^{(\mathbf{x})}$ instead of random Gaussian. Consequently, the learning objective function and the inference process are reformulated as follows.
$\begin{gathered}\mathcal{L}=\mathbb{E}_{(\mathbf{z}^{(\mathbf{x})},\mathbf{z}^{(\mathbf{y})}),t\sim\mathcal{U}(T)}\left\|(\sqrt{\bar{\alpha}_{t}}\mathbf{z}^{(\mathbf{x})}-\sqrt{1-\bar{\alpha}_{t}}\mathbf{z}^{(\mathbf{y})})-\bm{v}_{\theta}(\mathbf{z}_{t},t)\right\|^{2}_{2}.\end{gathered}$ (5)
$\begin{gathered}\hat{\mathbf{z}}^{(\mathbf{y})}_{T}=\mathbf{z}^{(\mathbf{x})},~{}~{}t=[T,...,1],\\ \hat{\mathbf{z}}^{(\mathbf{y})}_{t\rightarrow 0}=\sqrt{\bar{\alpha}_{t}}\cdot\hat{\mathbf{z}}^{(\mathbf{y})}_{t}-\sqrt{1-\bar{\alpha}_{t}}\cdot\bm{v}_{\theta}(\hat{\mathbf{z}}^{(\mathbf{y})}_{t},t),\hat{\mathbf{z}}^{(\mathbf{x})}_{t}=\sqrt{\bar{\alpha}_{t}}\cdot\bm{v}_{\theta}(\hat{\mathbf{z}}^{(\mathbf{y})}_{t},t)-\sqrt{1-\bar{\alpha}_{t}}\cdot\hat{\mathbf{z}}^{(\mathbf{y})}_{t},\\ \hat{\mathbf{z}}^{(\mathbf{y})}_{t-1}=\sqrt{\bar{\alpha}_{t-1}}\cdot\hat{\mathbf{z}}^{(\mathbf{y})}_{t\rightarrow 0}+\sqrt{1-\bar{\alpha}_{t-1}}\cdot\hat{\mathbf{z}}^{(\mathbf{x})}_{t},~{}~{}\hat{\mathbf{y}}=\mathcal{D}(\hat{\mathbf{z}}^{(\mathbf{y})}_{1\rightarrow 0}).\end{gathered}$ (6)
where $\hat{\mathbf{z}}^{(\mathbf{x})}_{t}$ denotes the predicted RGB noise of timestep $t$.
A.3 GenPercept: Deterministic One-step Perception
In the main paper, we suppose that blending with the RGB image is much simpler compared to Gaussian noise, and we increase the blending proportion of the RGB image to make it more difficult. This is realized by increasing the initial value $\beta_{0}$ and final values $\beta_{T}$ of the DDPM schedule. When all the beta values $\{\beta_{t}\}_{t=1}^{T}$ are increased to 1, and $\bar{\alpha}_{t}=\prod_{s=1}^{t}{(1-\beta_{s})}=0$, the formulation of eq. 4 to eq. 6 can be greatly simplified as follows.
$\begin{gathered}\mathbf{z}_{t}=\mathbf{z}^{(\mathbf{y})}_{t}=\mathbf{z}^{(\mathbf{x})},~{}~{}\mathcal{L}=\mathbb{E}_{(\mathbf{z}^{(\mathbf{x})},\mathbf{z}^{(\mathbf{y})}),t\sim\mathcal{U}(T)}\left\|-\mathbf{z}^{(\mathbf{y})}-\bm{v}_{\theta}(\mathbf{z}_{t},t)\right\|^{2}_{2},\\ \hat{\mathbf{z}}^{(\mathbf{y})}_{T}=\mathbf{z}^{(\mathbf{x})},~{}~{}\hat{\mathbf{z}}^{(\mathbf{y})}_{t-1}=-\hat{\mathbf{z}}^{(\mathbf{y})}_{t},~{}~{}\hat{\mathbf{z}}^{(\mathbf{y})}_{1\rightarrow 0}=-\bm{v}_{\theta}(\hat{\mathbf{z}}^{(\mathbf{y})}_{1},t=1),~{}~{}\hat{\mathbf{y}}=\mathcal{D}(\hat{\mathbf{z}}^{(\mathbf{y})}_{1\rightarrow 0}).\end{gathered}$ (7)
One-step prediction. The input of the U-Net is an RGB latent code, and the output becomes the negative value of the ground-truth latent code, with no relationship to the timestep $t$. Therefore, we set the number of timesteps $T$ to 1 while preserving the same performance.
$\begin{gathered}\mathcal{L}=\mathbb{E}_{(\mathbf{z}^{(\mathbf{x})},\mathbf{z}^{(\mathbf{y})})}\left\|-\mathbf{z}^{(\mathbf{y})}-\bm{v}_{\theta}(\mathbf{z}^{(\mathbf{x})},t=1)\right\|^{2}_{2},\\ \hat{\mathbf{z}}^{(\mathbf{y})}_{1\rightarrow 0}=-\bm{v}_{\theta}(\mathbf{z}^{(\mathbf{x})},t=1),~{}~{}\hat{\mathbf{y}}=\mathcal{D}(\hat{\mathbf{z}}^{(\mathbf{y})}_{1\rightarrow 0}).\end{gathered}$ (8)
Appendix B More Visualization Analysis
Quantitative Comparisons of Ablations. Visualization of the ablation study experiments in the main paper §2 is shown in fig. 7. The estimated depth detail remains comparable with a customized “DPT head”. Models trained with lower-quality data like “Taskonomy + Cityscapes” or without the pre-trained VAE decoder parameters suffer from a decline in the ability to predict details. We attribute them to the low-quality annotation and the oversized decoder, respectively.
Forward Diffusion Process. More detailed visualization of different noise forms and proportions in the forward diffusion process is shown in fig. 8. With larger $\beta$ values of the DDIM scheduler, the proportion of noise will be much higher during the forward diffusion process. When $\beta_{start},\beta_{end}$ reaches (1.0, 1.0), the noisy latent will be pure noise latent, which is Gaussian noise and RGB latent for (a) and (b), respectively.
Appendix C An Extra Attempt on Human Pose Estimation
Task Definition Human Pose Estimation is a task aimed at determining the spatial configuration of a person or object in a given image or video. This involves identifying and predicting the coordinates of particular keypoints.
Implementation Details For human keypoint detection, we use Simple Baseline (Xiao et al., 2018a) for person detection and conduct training on the COCO training set with 15K training samples. Performance is evaluated on the COCO (Lin et al., 2014) validation set. As shown in fig. 9, it is generalizable to unseen objects in the training set. To evaluate the performance on COCO, we use the customized keypoint head of ViTPose (Xu et al., 2022) for decoding the output. The quantitative results compared with a generalist model, Painter(Wang et al., 2023) is shown in table 11.
Table 11: Pose estimation on COCO.
Metrics AP $\uparrow$ AP .5 $\uparrow$ AP .75 $\uparrow$ AP (M) $\uparrow$ AP (L) $\uparrow$
Painter (Wang et al., 2023) 0.721 0.900 0.781 0.686 0.786
GenPercept 0.752 0.907 0.824 0.691 0.778
Figure 9: Generalized test results of keypoint detection.
Appendix D Implementation Details
Implementation Details of Preliminary Section in the Main Paper. Unless specified otherwise, we train for 30000 iterations to estimate the affine-invariant monocular depth on a mixture of the Hypersim (Roberts et al., 2021) (90%) and Virtual KITTI (Cabon et al., 2020) (10%) datasets. We freeze the VAE AutoEncoder and fine-tune the U-Net of Stable Diffusion v2.1 to estimate the ground-truth label latent, with a resolution of (480, 640), a batch size of 32, and a learning rate of 3e-5. The multi-resolution noise (Ke et al., 2024) is employed for Gaussian noise and not used for RGB noise (Lee et al., 2024) by default. For inference, the ensemble size and denoising steps are set to 1 and 10, respectively. Evaluation results of absolute relative error (AbsRel) and $\delta_{1}$ are reported on five unseen monocular depth datasets, including KITTI (Geiger et al., 2013), NYU (Silberman et al., 2012), ScanNet (Dai et al., 2017), DIODE (Vasiljevic et al., 2019), and ETH3D (Schops et al., 2017).
Appendix E More Quantitative and Qualitive Experiments
E.1 Monocular Depth Estimation
More Qualitative Evaluation. We show the robustness of the monocular depth estimation model in diverse scenes in Fig. 11. Compared to DPT (Ranftl et al., 2021), our Genpercept performs better on estimating details and shows much better robustness even on some sketches. Compared to Marigold (Ke et al., 2024), our method achieves better relative depth visualization.
E.2 Surface Normal Estimation
More Qualitative Evaluation. In Fig. 12, we showcase more qualitative results for the surface normal estimation. DSINE (Bae & Davison, 2024b) is trained on 160K images of 10 datasets, including both real and synthetic data. Our GenPercept is only trained on one synthetic dataset (Hypersim (Roberts et al., 2021)), and can estimate much more detailed surface normal maps.
E.3 Dichotomous Image Segmentation
More Quantitative Evaluation. To conduct a comprehensive evaluation, we compare our approach with numerous previous models including models for medical image segmentation (Ronneberger et al., 2015), salient object detection (Qin et al., 2019; Zhao et al., 2020; Wei et al., 2020; Chen et al., 2020; Qin et al., 2020), camouflaged object detection (Fan et al., 2021a; Mei et al., 2021), semantic segmentation (Zhao et al., 2017; Chen et al., 2018a; Wang et al., 2020; Yu et al., 2018; Zhao et al., 2018; Howard et al., 2019; Fan et al., 2021c; Nirkin et al., 2021) and the model IS-Net (Qin et al., 2022) specifically trained for DIS. As shown in table 12, our proposed model significantly outperforms other methods across most evaluation metrics on this challenging dataset. The results highlight the effectiveness of our approach for DIS.
Table 12: Quantitative results on DIS5K validation and testing sets.
Dataset Metric
UNet
BASNet
GateNet
F3Net
GCPANet
U2Net
SINetV2
PFNet
PSPNet
DLV3+
HRNet
BSV1
ICNet
MBV3
STDC
HySM
IS-Net
Ours
DIS-VD $maxF_{\beta}\uparrow$ 0.692 0.731 0.678 0.685 0.648 0.748 0.665 0.691 0.691 0.660 0.726 0.662 0.697 0.714 0.696 0.734 0.791 0.857
$F^{w}_{\beta}\uparrow$ 0.586 0.641 0.574 0.595 0.542 0.656 0.584 0.604 0.603 0.568 0.641 0.548 0.609 0.642 0.613 0.640 0.717 0.835
${}M~{}\downarrow$ 0.113 0.094 0.110 0.107 0.118 0.090 0.110 0.106 0.102 0.114 0.095 0.116 0.102 0.092 0.103 0.096 0.074 0.040
$S_{\alpha}\uparrow$ 0.745 0.768 0.723 0.733 0.718 0.781 0.727 0.740 0.744 0.716 0.767 0.728 0.747 0.758 0.740 0.773 0.813 0.870
$E_{\phi}^{m}\uparrow$ 0.785 0.816 0.783 0.800 0.765 0.823 0.798 0.811 0.802 0.796 0.824 0.767 0.811 0.841 0.817 0.814 0.856 0.934
${\rm HCE}_{\gamma}\downarrow$ 1337 1402 1493 1567 1555 1413 1568 1606 1588 1520 1560 1660 1503 1625 1598 1324 1116 1511
DIS-TE1 $maxF_{\beta}\uparrow$ 0.625 0.688 0.620 0.640 0.598 0.694 0.644 0.646 0.645 0.601 0.668 0.595 0.631 0.669 0.648 0.695 0.740 0.841
$F^{w}_{\beta}\uparrow$ 0.514 0.595 0.517 0.549 0.495 0.601 0.558 0.552 0.557 0.506 0.579 0.474 0.535 0.595 0.562 0.597 0.662 0.814
${}M~{}\downarrow$ 0.106 0.084 0.099 0.095 0.103 0.083 0.094 0.094 0.089 0.102 0.088 0.108 0.095 0.083 0.090 0.082 0.074 0.038
$S_{\alpha}\uparrow$ 0.716 0.754 0.701 0.721 0.705 0.760 0.727 0.722 0.725 0.694 0.742 0.695 0.716 0.740 0.723 0.761 0.787 0.868
$E_{\phi}^{m}\uparrow$ 0.750 0.801 0.766 0.783 0.750 0.801 0.791 0.786 0.791 0.772 0.797 0.741 0.784 0.818 0.798 0.803 0.820 0.918
${\rm HCE}_{\gamma}\downarrow$ 233 220 230 244 271 224 274 253 267 234 262 288 234 274 249 205 149 204
DIS-TE2 $maxF_{\beta}\uparrow$ 0.703 0.755 0.702 0.712 0.673 0.756 0.700 0.720 0.724 0.681 0.747 0.680 0.716 0.743 0.720 0.759 0.799 0.876
$F^{w}_{\beta}\uparrow$ 0.597 0.668 0.598 0.620 0.570 0.668 0.618 0.633 0.636 0.587 0.664 0.564 0.627 0.672 0.636 0.667 0.728 0.852
${}M~{}\downarrow$ 0.107 0.084 0.102 0.097 0.109 0.085 0.099 0.096 0.092 0.105 0.087 0.111 0.095 0.083 0.092 0.085 0.070 0.035
$S_{\alpha}\uparrow$ 0.755 0.786 0.737 0.755 0.735 0.788 0.753 0.761 0.763 0.729 0.784 0.740 0.759 0.777 0.759 0.794 0.823 0.884
$E_{\phi}^{m}\uparrow$ 0.796 0.836 0.804 0.820 0.786 0.833 0.823 0.829 0.828 0.813 0.840 0.781 0.826 0.856 0.834 0.832 0.858 0.938
${\rm HCE}_{\gamma}\downarrow$ 474 480 501 542 574 490 593 567 586 516 555 621 512 600 556 451 340 480
DIS-TE3 $maxF_{\beta}\uparrow$ 0.748 0.785 0.726 0.743 0.699 0.798 0.730 0.751 0.747 0.717 0.784 0.710 0.752 0.772 0.745 0.792 0.830 0.885
$F^{w}_{\beta}\uparrow$ 0.644 0.696 0.620 0.656 0.590 0.707 0.641 0.664 0.657 0.623 0.700 0.595 0.664 0.702 0.662 0.701 0.758 0.862
${}M~{}\downarrow$ 0.098 0.083 0.103 0.092 0.109 0.079 0.096 0.092 0.092 0.102 0.080 0.109 0.091 0.078 0.090 0.079 0.064 0.035
$S_{\alpha}\uparrow$ 0.780 0.798 0.747 0.773 0.748 0.809 0.766 0.777 0.774 0.749 0.805 0.757 0.780 0.794 0.771 0.811 0.836 0.883
$E_{\phi}^{m}\uparrow$ 0.827 0.856 0.815 0.848 0.801 0.858 0.849 0.854 0.843 0.833 0.869 0.801 0.852 0.880 0.855 0.857 0.883 0.951
${\rm HCE}_{\gamma}\downarrow$ 883 948 972 1059 1058 965 1096 1082 $1$ 999 1049 1146 1001 1136 1081 887 687 973
DIS-TE4 $maxF_{\beta}\uparrow$ 0.759 0.780 0.729 0.721 0.670 0.795 0.699 0.731 0.725 0.715 0.772 0.710 0.749 0.736 0.731 0.782 0.827 0.848
$F^{w}_{\beta}\uparrow$ 0.659 0.693 0.625 0.633 0.559 0.705 0.616 0.647 0.630 0.621 0.687 0.598 0.663 0.664 0.652 0.693 0.753 0.829
${}M~{}\downarrow$ 0.102 0.091 0.109 0.107 0.127 0.087 0.113 0.107 0.107 0.111 0.092 0.114 0.099 0.098 0.102 0.091 0.072 0.049
$S_{\alpha}\uparrow$ 0.784 0.794 0.743 0.752 0.723 0.807 0.744 0.763 0.758 0.744 0.792 0.755 0.776 0.770 0.762 0.802 0.830 0.854
$E_{\phi}^{m}\uparrow$ 0.821 0.848 0.803 0.825 0.767 0.847 0.824 0.838 0.815 0.820 0.854 0.788 0.837 0.848 0.841 0.842 0.870 0.938
${\rm HCE}_{\gamma}\downarrow$ 3218 3601 3654 3760 3678 3653 3683 3803 3806 3709 3864 3999 3690 3817 3819 3331 2888 3799
Overall DIS-TE (1-4) $maxF_{\beta}\uparrow$ 0.708 0.752 0.694 0.704 0.660 0.761 0.693 0.712 0.710 0.678 0.743 0.674 0.711 0.729 0.710 0.757 0.799 0.863
$F^{w}_{\beta}\uparrow$ 0.603 0.663 0.590 0.614 0.554 0.670 0.608 0.624 0.620 0.584 0.658 0.558 0.622 0.658 0.628 0.665 0.726 0.839
${}M~{}\downarrow$ 0.103 0.086 0.103 0.098 0.112 0.083 0.101 0.097 0.095 0.105 0.087 0.110 0.095 0.085 0.094 0.084 0.070 0.039
$S_{\alpha}\uparrow$ 0.759 0.783 0.732 0.750 0.728 0.791 0.747 0.756 0.755 0.729 0.781 0.737 0.758 0.770 0.754 0.792 0.819 0.872
$E_{\phi}^{m}\uparrow$ 0.798 0.835 0.797 0.819 0.776 0.835 0.822 0.827 0.819 0.810 0.840 0.778 0.825 0.850 0.832 0.834 0.858 0.936
${\rm HCE}_{\gamma}\downarrow$ 1202 1313 1339 1401 1395 1333 1411 1427 1442 1365 1432 1513 1359 1457 1426 1218 1016 1364
More Qualitative Evaluation. As shown in Fig. 13, our method yields more refined segmentation results, providing cleaner foreground masks. It also produces precise outputs for intricate lines that are challenging for previous approaches to handle.
Cross Dataset Evaluation. To test the generalization ability of our model, we randomly select some images from other datasets(Agustsson & Timofte, 2017; Lin et al., 2014; Shao et al., 2019) for experiments. As shown in Fig. 10, compared to IS-Net and IS-Net-General-UseQin et al. (2022), our approach exhibits finer segmentation quality across diverse images, providing cleaner foreground masks. It is noteworthy that IS-Net-General-Use is fine-tuned on extra datasets to enhance generalization, which indicates that our method has a stronger generalization ability.
Figure 10: Cross dataset comparison of our model and other models.
E.4 Image Matting
More Qualitative Evaluation. In Fig. 14, we showcase more qualitative results for the image matting task. It is worth noting that our model works well in various resolutions, light environments, human poses, and human orientations. Please zoom in for better visualization and more details.
E.5 Image Segmentation
More Qualitative Evaluation. In Fig. 15, we showcase more qualitative results for the image segmentation task. Our method shows much robustness on the trained categories of complex in-the-wild images. Due to the limited annotation categories and little negative label of “unknown category”, it sometimes fails in outdoor scenes and unseen categories such as cats and cars. Please zoom in for better visualization and more details.
E.6 Human Pose Estimation
More Qualitative Evaluation. In Fig.16, we showcase more qualitative results for the human pose estimation. We conduct experiments on MHP datasetLi et al. (2017), and we use mmposeContributors (2020) to render the human pose following the setting of Bai et al. (2023).
Figure 11: More qualitative results for monocular depth estimation.
Figure 12: More qualitative comparisons of surface normal estimation. Our GenPercept can achieve more detailed results, even compared to the competitive CVPR2024 method DSINE Bae & Davison (2024b). Note that these two visualization coordinate systems are slightly different.
Figure 13: Qualitative results of dichotomous image segmentation.
Figure 14: More qualitative results for image matting.
Figure 15: More qualitative results for image matting.
Figure 16: More qualitative results for human pose estimation. (Left: original Image, Mid: prediction, Right: ground truth)

### v4

@@
 Extensive pre-training with large data is indispensable for downstream geometry and semantic visual perception tasks.
-Thanks to large-scale text-to-image (T2I) pretraining, recent works show promising results by simply fine-tuning T2I diffusion models for dense perception tasks.
-However, several crucial design decisions in this process still lack comprehensive justification, encompassing the necessity of the multi-step stochastic diffusion mechanism, training strategy, inference ensemble strategy, and fine-tuning data quality.
+Thanks to large-scale text-to-image (T2I) pretraining, recent works show promising results by simply fine-tuning T2I diffusion models for a few dense perception tasks.
+However, several crucial design decisions in this process still lack comprehensive justification, encompassing the necessity of the multi-step diffusion mechanism, training strategy, inference ensemble strategy, and fine-tuning data quality.
 In this work, we conduct a thorough investigation into critical factors that affect transfer efficiency and performance when using diffusion priors.
 Our key findings are: 1) High-quality fine-tuning data is paramount for both semantic and geometry perception tasks.
-2) The stochastic nature of diffusion models has a slightly negative impact on deterministic visual perception tasks.
-3) Apart from fine-tuning the diffusion model with only latent space supervision, task-specific image-level supervision is beneficial to enhance fine-grained details.
-These observations culminate in the development of GenPercept, an effective deterministic one-step fine-tuning paradigm tailed for dense visual perception tasks.
-Different from the previous multi-step methods, our paradigm has a much faster inference speed, and can be seamlessly integrated with customized perception decoders and loss functions for image-level supervision, which is critical to improving the fine-grained details of predictions.
-Comprehensive experiments on diverse dense visual perceptual tasks, including monocular depth estimation, surface normal estimation, image segmentation, and matting, are performed to demonstrate the remarkable adaptability and effectiveness of our proposed method.
-Code is at:
-https://github.com/aim-uofa/GenPercept/
+2) As a special case of the diffusion scheduler by setting its hyper-parameters, the multi-step generation can be simplified to a one-step fine-tuning paradigm without any loss of performance.
+3) Apart from fine-tuning the diffusion model with only latent space supervision, task-specific supervision is beneficial to enhance fine-grained details.
+These observations culminate in the development of GenPercept, an effective deterministic one-step fine-tuning paradigm tailored for dense visual perception tasks exploiting diffusion priors.
+Different from the previous multi-step methods, our paradigm offers a much faster inference speed, and can be seamlessly integrated with customized perception decoders and loss functions for task-specific supervision, which can be critical for improving the fine-grained details of predictions.
+Comprehensive experiments on a diverse set of dense visual perceptual tasks, including monocular depth estimation, surface normal estimation, image segmentation, and matting, are performed to demonstrate the remarkable adaptability and effectiveness of our proposed method.
 1 Introduction
 Recent studies have explored the transferability of text-to-image (T2I) diffusion models to dense visual perception tasks, such as geometry estimation (Ke et al., 2024; Lee et al., 2024; Fu et al., 2024b; Gui et al., 2024; Ye et al., 2024), image segmentation (Van Gansbeke & De Brabandere, 2024; Lee et al., 2024), and inverse rendering (Chen et al., 2024; Kocsis et al., 2024; Zeng et al., 2024).
-While these works have demonstrated impressive results by repurposing diffusion models for estimating fine-grained geometric and semantic dense prediction maps, the critical design choices made in transferring diffusion models to other dense perception tasks still lack comprehensive justification.
-This makes it challenging to determine the optimal strategy for achieving the best performance.
+While these works have demonstrated impressive results by repurposing diffusion models for estimating geometric and semantic dense prediction maps, the critical design choices made in transferring diffusion models to other dense perception tasks still lack comprehensive justification.
+This makes it challenging to determine the optimal strategy for achieving optimal performance.
 For example, Ke et al.
 (2024) align the visual perception process with the denoising process of Stable Diffusion by fine-tuning all U-Net parameters.
-They highlight the significance of ”multi-resolution noise” in the forward diffusion process during training, aiming to obtain clean predictions by gradually removing Gaussian noise.
-On the other hand, (Lee et al., 2024) modify the forward diffusion process by interpolating perception annotations with RGB images instead of using Gaussian noise, and only train the low-rank adaptation (LoRA) Hu et al.
-(2022) parameters while keeping the U-Net frozen.
-To the best of our knowledge, the effective components of these approaches have not been thoroughly investigated, leading to uncertainty about which aspects are most crucial for success.
-In this work, we examine the design space of repurposing diffusion models for dense visual perception tasks, and explore the key question: What factors are important when adapting diffusion models for general dense perception tasks?
+They highlight the significance of “multi-resolution noise” in the forward diffusion process during training, aiming to obtain clean predictions by gradually removing Gaussian noise.
+On the other hand, Lee et al.
+(2024) modify the forward diffusion process by interpolating perception annotations with RGB images instead of using Gaussian noise, and only train the low-rank adaptation (LoRA) (Hu et al., 2022) parameters while keeping the U-Net frozen.
+To our knowledge, the effective components of these approaches have not been thoroughly investigated, and it is unclear which design choices contribute most to the success.
+In this work, we examine the design space of repurposing diffusion models for dense visual perception tasks, and attempt to answer the key question: What are the important design choices when adapting diffusion models for general dense perception tasks?
 To answer this question, we rethink the importance of both fine-tuning protocols and fine-tuning data.
 From the perspective of fine-tuning protocols, we categorize recent methods into two main groups: stochastic multi-step generation and deterministic multi-step generation.
-We explore several critical design spaces, including the diffusion mechanism, key architectural components, training methodologies, and data quality.
-Our key observations are as follows: 1) The stochastic nature of diffusion models somewhat conflicts with the deterministic requirements of perception tasks, and the amount of noise used significantly impacts perception performance.
-2) Strict adherence to traditional diffusion processes is unnecessary.
+We explore several critical design choices, including the diffusion mechanism, key architectural components, training methodologies, and data quality.
+Our key observations are as follows: 1) By setting the hyperparameters of the diffusion scheduler to particular values, the multi-step generation can be simplified to a one-step fine-tuning paradigm without any loss of performance.
+2) Strict adherence to traditional diffusion processes appears to be unnecessary.
 Single-step inference provides similar performance with significantly faster execution.
-3) High-quality synthetic fine-tuning data is crucial for geometry perception tasks.
+3) High-quality synthetic fine-tuning data is crucial for several perception tasks.
 From the perspective of fine-tuning data quality, we conduct comprehensive dataset ablation studies on both synthetic datasets and real-world datasets.
 Based on the aforementioned observations, we propose GenPercept (see Fig.
-1), a deterministic fine-tuning paradigm featuring a remarkably simple one-step inference pipeline, an optional customized decoder, and an easily adaptable pixel-specific customized loss.
-We conduct extensive quantitative and qualitative experiments on a wide range of fundamental visual dense perception tasks, including monocular depth estimation, surface normal estimation, image segmentation, and matting to demonstrate the effectiveness and universality of our method.
-In conclusion, our contributions can be summarized as follows: 1) We systematically analyze the design space of current fine-tuning protocols, considering both model architecture and dataset selection, through comprehensive ablation studies.
+1), a deterministic fine-tuning paradigm featuring a remarkably simple one-step inference pipeline, an optional customized decoder, and an easily adaptable pixel-wise loss.
+We conduct extensive quantitative and qualitative experiments on a wide range of visual dense perception tasks, including monocular depth estimation, surface normal estimation, image segmentation, and matting to demonstrate the effectiveness and generalization capability of our method.
+In conclusion, our contributions can be summarized as follows: 1) We systematically analyze the design space of fine-tuning protocols, considering both model architecture and dataset selection, through comprehensive ablation studies.
 2) Based on these insights, we propose GenPercept, a simple paradigm that harnesses the power of the pre-trained UNet from diffusion models for generalizable dense visual perception tasks.
@@
 2 preliminary
-Standard diffusion models (Rombach et al., 2022; Chen et al., 2023; Song et al., 2020; Ho et al., 2020) define a forward process to inject Gaussian noise $\bm{\epsilon}$ into the input data and a reverse denoising process to estimate noises and generate clean images with a denoiser $\bm{v}_{\theta}$.
-While diffusion models generate images from noise, the visual perception tasks take RGB images $\mathbf{x}$ as input and estimate the visual perception labels $\mathbf{y}$.
-Aiming at transferring the prior knowledge of pre-trained diffusion models to computer vision tasks, previous works have proposed to align with the diffusion process and reformulate visual perception tasks as a multi-step denoising process, especially on the monocular depth estimation task.
-We review the main high-level ideas of previous algorithms in this section and offer detailed formulations in the supplementary material.
-Recent works mainly explore two alignment choices.
-For the stochastic multi-step generation (Ke et al., 2024; Fu et al., 2024b; Gui et al., 2024), the perception tasks are reframed as an RGB image-conditioned denoising process.
-For training, they add random Gaussian noise $\bm{\epsilon}$ to the ground truth latent $\mathbf{z}^{(\mathbf{y})}$ in the forward diffusion process.
-For inference, the RGB latent $\mathbf{z}^{(\mathbf{x})}$ is concatenated with the noisy depth latent $\hat{\mathbf{z}}^{(\mathbf{y})}_{t}$ before being sent to U-Net $\bm{v}_{\theta}$ for the $t$-th denoising timestep.
-Clean depth latent $\hat{\mathbf{z}}^{(\mathbf{y})}_{0}$ will be available after denoising T times.
-The inference denoising process and forward diffusion process are shown in fig.
-1 (a) and fig.
-2 (a) separately.
-These methods perform well but face two significant drawbacks.
-Firstly, its stochastic nature conflicts with the deterministic nature of perception tasks, which may have a negative impact on performance.
-Secondly, the high-computation ensemble strategy poses challenges across various perception tasks.
-To integrate the deterministic requirements of visual perception tasks with the stochastic diffusion models, deterministic multi-step generation methods such as DMP (Lee et al., 2024) reformulate the forward diffusion process as a blending interpolation between RGB images and ground-truth labels.
-The inference denoising process and training forward diffusion process are shown in fig.
-1 (b) and fig.
-2 (a), respectively.
-Although it addresses the uncertainty inherent in the process, the persistence of RGB texture in the generated perception predictions can detrimentally impact accuracy when visualizing the estimated outcomes.
-This phenomenon is particularly evident in tasks such as surface normal estimation, as exemplified in fig.
-2 (c).
-Figure 2: Visualization of different noise forms and proportions in the forward diffusion process.
-(a) Compared to Gaussian Noise, interpolating RGB images with ground-truth labels can bring less image nature loss.
-(b) The relationship between the proportion of noise $\sqrt{\bar{\alpha}_{t}}$ and the ($\beta_{start}$, $\beta_{end}$) hyperparameters.
-(Please see the supplementary for detailed formulations.) (c) Low $\beta$ values during the training of deterministic multi-step generation will lead to the retention of RGB information.
-Increasing $\beta$ enhances the training difficulty and addresses this issue.
+We take the latent diffusion model as an example.
+To model the data distribution, the idea of the diffusion model (Rombach et al., 2022; Chen et al., 2023; Song et al., 2020; Ho et al., 2020) is to randomly sample a noise $\mathbf{z}^{(\mathbf{y})}_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ and sequentially denoise it into a $\mathbf{z}^{(\mathbf{y})}_{0}$, which is distributed according to the data.
+In the forward diffusion process, $\mathbf{z}^{(\mathbf{y})}_{t}$ is sampled by $\mathbf{z}^{(\mathbf{y})}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{z}^{(\mathbf{y})}+\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon}$, where $\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, and $\bar{\alpha}_{t}=\prod_{s=1}^{t}{(1-\beta_{s})}$.
+The variance schedule $\{\beta_{t}\in(0,1)\}_{t=1}^{T}$ is interpolated between $\beta_{start}$ and $\beta_{end}$ with $T$ steps, where larger values of $(\beta_{start},\beta_{end})$ correspond to smaller $\bar{\alpha}_{t}$ values, i.e., smaller proportion of noise.
+In the reverse process, Salimans & Ho (2021) use a “v-prediction” objective, where a denoiser $\bm{v}_{\theta}$ minimizes the following:
+$\begin{gathered}\mathcal{L}=\mathbb{E}_{\mathbf{z}^{(\mathbf{y})},\bm{\epsilon}\sim\mathcal{N}(0,I),t\sim\mathcal{U}(T)}\left\|(\sqrt{\bar{\alpha}_{t}}\bm{\epsilon}-\sqrt{1-\bar{\alpha}_{t}}\mathbf{z}^{(\mathbf{y})})-\bm{v}_{\theta}(\mathbf{z}_{t},t)\right\|^{2}_{2}.\end{gathered}$ (1)
+To fully leverage the pre-trained prior of the diffusion models for dense prediction tasks, previous works have reformulated these tasks as a multi-step denoising process, especially on monocular depth estimation.
+Given a data pair ($\mathbf{z}^{(\mathbf{x})}$, $\mathbf{z}^{(\mathbf{y})}$) where $\mathbf{z}^{(\mathbf{x})}$ is the observation and $\mathbf{z}^{(\mathbf{y})}$ is the prediction target, stochastic multi-step generation methods (Ke et al., 2024; Fu et al., 2024b; Gui et al., 2024) such as Marigold (Ke et al., 2024) add $\mathbf{z}^{(\mathbf{x})}$ as an additional input to the denoiser $\bm{v}_{\theta}$, and use $\bm{v}_{\theta}(\mathbf{z}^{(\mathbf{y})}_{t},\mathbf{z}^{(\mathbf{x})},t)$ to predict $\mathbf{z}^{(\mathbf{y})}$.
+By contrast, deterministic multi-step generation methods such as DMP (Lee et al., 2024) take the observation $\mathbf{z}^{(\mathbf{x})}$ as a deterministic noise and compose $\mathbf{z}^{(\mathbf{y})}_{t}$ as a blend between $\mathbf{z}^{(\mathbf{x})}$ and $\mathbf{z}^{(\mathbf{y})}$:
+$\begin{gathered}\mathbf{z}_{t}=\mathbf{z}^{(\mathbf{y})}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{z}^{(\mathbf{y})}+\sqrt{1-\bar{\alpha}_{t}}\mathbf{z}^{(\mathbf{x})},~{}~{}t=[1,...,T],\\ \end{gathered}$ (2)
+The denoising process and forward diffusion process of these two categories are illustrated in Fig. 1 and Fig. 2 (a). We offer detailed formulations in the supplementary material.
+Figure 2: Illustration of different noise forms and proportions in the forward diffusion process.
+(a) Visualization of interpolating ground-truth labels with Gaussian noise and RGB noise.
+(b) The relationship between the noise proportion $\sqrt{\bar{\alpha}_{t}}$ and the ($\beta_{start}$, $\beta_{end}$) hyperparameters.
+(c) Small $(\beta_{start},\beta_{end})$ values during the training of deterministic multi-step generation tend to lead to unclean estimation, which contains some RGB information.
+Enlarging them may alleviate this issue.
 3 Diffusion Models for Visual Perception Tasks
-In this section, we explore the necessity and highlight the findings of the multi-step stochastic diffusion mechanism, prior knowledge hides behind the architectural components, training strategy, and fine-tuning data quality.
-The default experimental setting is similar to Ke et al.
-(2024) and can be found in supplementary material.
-We select stochastic method Marigold (Ke et al., 2024) and the deterministic method DMP (Lee et al., 2024) as our baseline.
+In this section, we explore the necessity and highlight the findings of the multi-step diffusion mechanism, the architectural components, training strategy, and fine-tuning data quality.
+We select the stochastic method Marigold (Ke et al., 2024) and the deterministic method DMP (Lee et al., 2024) as our baseline methods.
+The default experimental setting here is similar to Ke et al.
+(2024) and can be found in the supplementary material.
 3.1 The Form and Proportion of Noise in the Forward Diffusion Process
-In our initial attempt to integrate the core concepts of Marigold and DMP within a unified codebase, we observed that the performance of the ”baseline DMP” was significantly inferior to that of the ”baseline Marigold.” This discrepancy led us to uncover two notable phenomena: 1) Although deterministic methods are theoretically expected to outperform stochastic, multi-step generation processes for visual perception tasks due to their reduced dependency on random Gaussian noise, our findings contradicted this assumption, with deterministic methods showing inferior performance.
-2) The ”multi-resolution noise” strategy introduced by Marigold substantially enhances accuracy.
-To further investigate these phenomena, we hypothesize that the challenge associated with denoising from Gaussian noisy latents is substantially more difficult than that from RGB-blended noisy latents.
-These two different blending paradigms are illustrated in fig.
-2(a).
-We argue that the Gaussian noisy latent less closely approximates natural images and potentially increases the complexity of training, which can enhance the training difficulty.
-To prove it, we modified the noise blending proportions for both methods by adjusting the $\beta$ values within the DDPM scheduler, effectively increasing the noise component during the forward diffusion process, as depicted in fig.
-2(b).
-Our quantitative and qualitative analyses, presented in table 1 and fig.
-2(c), indicate that increasing training challenges lead to improved model performance.
-As the $\beta$ value increases, the impact of ”multi-resolution noise” diminishes.
-However, unlike DMP, which continues to show performance improvements, Marigold’s performance begins to be slightly unstable but shows a rough improvement trend when the $\beta$ value is sufficiently high.
-We consider it brought by the randomness of the Gaussian Noise, which dominates most components of the input latent.
-When $\beta$ reaches 1, the noisy label latent input effectively becomes a pure noise latent, akin to the RGB latent in DMP.
-Consequently, the target output of the ”v-prediction” diffusion model shifts to the negative value of the ground truth latent, eliminating any intermediate state of blending between RGB noise latent and ground-truth latent.
-The denoising process becomes repeats of “predict target label $\hat{\mathbf{z}}^{(\mathbf{y})}_{0}$ from RGB image $\mathbf{z}^{(\mathbf{x})}$ and add (RGB) noise back to the RGB image $\mathbf{z}^{(\mathbf{x})}$”, whose estimation results in any timestep $t$ are exactly identical.
-Therefore, we propose to reduce the DDIM steps to one and call it ”deterministic one-step perception”.
-The resulting inference speeds are significantly faster, with performance remaining largely consistent.
-We consider it as “our baseline” for the subsequent analyses.
+For the training process of Marigold and DMP, the timestep $t$ is sampled to control the proportion of noise added to the ground truth, and the network is trained to recover a clean ground truth from a noisy latent.
+For smaller timesteps such as “$t=200$”, as illustrated in Fig.
+2(a), the input to the network retains significant ground truth information (e.g., the purple color of the normal).
+Therefore, we hypothesize that a certain level of ground-truth label information being part of the input makes the network comparatively be easier to recover the clean ground truth latent than the case of the absence of any ground truth information during training.
+This can limit the network’s capacity to learn comprehensive knowledge and lead to unsatisfactory performance, as it is known that networks can become lazy that tend to exploit “shortcuts”—the ground truth labels in the input in our case.
+On the other hand, for inference there is no such ground truth available, causing disparity of input signals between inference and training.
+We note that T2I tasks may suffer less due to their stochastic nature, namely, converting a text prompt to a generated image is one-to-many mapping process.
+To alleviate this issue, we attempt to control the blending proportion by changing the ($\beta_{start}$, $\beta_{end}$) values of the diffusion model’s DDPM scheduler.
+As shown in Fig.
+2(b) and Fig.
+2 of the supplementary, training with a ($\beta_{start}$, $\beta_{end}$) value of (1.0, 1.0) achieves the best rank performance for both Gaussian noise and RGB noise, which is demonstrated in Table 1.
+Rather than achieving consistent performance improvement while increasing the noise proportion, we observed that Marigold’s performance begins to be slightly unstable when the ($\beta_{start}$, $\beta_{end}$) values are sufficiently high.
+Experiments of varying the random seed during both the training and inference process are conducted to rule out the influence of randomness.
+The results show that we may not be able to exactly find a set of unique values for ($\beta_{start}$, $\beta_{end}$) to achieve the best accuracy, as the final accuracy can be affected by a few other factors besides the aforementioned one.
+Additionally, when ($\beta_{start}$, $\beta_{end}$) are equal to 1, the noise proportion $\sqrt{\bar{\alpha}_{t}}$ is equal to 0, and the formulation of DMP can be derived from Eq.
+1 and Eq.
+8 as follows.
+$\begin{gathered}\mathbf{z}_{t}=\mathbf{z}^{(\mathbf{y})}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{z}^{(\mathbf{y})}+\sqrt{1-\bar{\alpha}_{t}}\mathbf{z}^{(\mathbf{x})}=\mathbf{z}^{(\mathbf{x})},~{}~{}t=[1,...,T],\\ \mathcal{L}=\mathbb{E}_{\mathbf{z}^{(\mathbf{y})},\bm{\epsilon}\sim\mathcal{N}(0,I),t\sim\mathcal{U}(T)}\left\|-\mathbf{z}^{(\mathbf{y})}-\bm{v}_{\theta}(\mathbf{z}_{t},t)\right\|^{2}_{2}.\end{gathered}$ (3)
+In this case, the output of the denoiser $\bm{v}_{\theta}(\cdot,\cdot)$ is enforced to learn the negative value of the ground truth latent for each step, and the multi-step denoising is equivalent to the single-step denoising.
+We propose to reduce the DDIM steps of DMP to one and call it “deterministic one-step perception”.
+The resulting inference can be significantly faster, with performance remaining almost unchanged.
+We name this as “our baseline” for the subsequent analysis.
 Table 1: Comprehensive quantitative comparisons about the impact of noise forms and proportions in the forward diffusion process on monocular depth estimation.
-Visualizations of different noise forms and the effect of $\beta$ values are shown in fig.
+Visualizations of different noise forms and the effect of ($\beta_{start}$, $\beta_{end}$) values are shown in Fig.
 2.
 The performance of DMP improves steadily, while Marigold shows initial improvements followed by a decline.
-When $\beta$ reaches 1, the inference process can be reduced to one step without compromising performance.
+When $\beta_{start}$ and $\beta_{end}$ are equal to 1, the inference process can be reduced to one step without compromising performance.
 “Rank” means the average rank of ten evaluation performance (smaller is better).
-Type Noise Multi-res Steps ($\beta_{start}$, $\beta_{end}$) KITTI NYU ScanNet DIODE ETH3D Rank
+[…]
 […]
@@
 Finding 1.
-The inherent stochasticity of diffusion models can lead to minor, unstable performance fluctuations in deterministic perception tasks.
-By increasing the noise proportion, the multi-step generation can be simplified to a one-step fine-tuning paradigm without any loss of performance in both stochastic and deterministic methods.
+By setting the ($\beta_{start}$, ${\beta_{end}}$) values to 1, the multi-step generation is simplified to a one-step fine-tuning paradigm without any loss of performance in both stochastic and deterministic methods, e.g., Marigold (Ke et al., 2024) and DMP (Lee et al., 2024) respectively.
 3.2 Where Does the Rich Visual Knowledge Reside in Diffusion Models?
@@
 Denoiser.
-We reinitialize the U-Net parameter and train it from scratch on the same datasets.
-Without prior knowledge of large data from LAION-5B, the network performs poorly and loses the generalization.
-This indicates that most of the prior knowledge is stored in the denoiser module.
+We reinitialize the U-Net parameters and train the network from scratch on the same datasets.
+Without prior knowledge of large data from LAION-5B, the network performs poorly and loses the generalization capability.
+This indicates that most of the prior knowledge is stored in the U-Net module.
 VAE AutoEncoder.
@@
 Customized Head and Loss.
-The deterministic one-step perception pipeline enables customized heads and loss functions for training.
-By utilizing a DPT decoder or the loss functions of DepthAnythingv2 (Yang et al., 2024b), we can implement a lightweight decoder that supervises pixel-wise information at a higher resolution rather than latent features at 1/8 resolution.
-This approach can not only accelerate inference times but also enhance the acquisition of fine-grained details.
+The deterministic one-step perception pipeline enables customized heads and loss functions.
+By utilizing a DPT decoder (Ranftl et al., 2021) and the loss functions of DepthAnythingv2 (Yang et al., 2024b), we can implement a lightweight decoder that supervises pixel-wise information at a higher resolution rather than latent features at 1///8 resolution.
+This approach can accelerate inference times and enhance the acquisition of fine-grained details.
 Finding 2.
-The primary perceptual prior knowledge of diffusion models is encapsulated within the denoiser. Customized heads and loss functions can achieve enhanced inference speed and fine-grained details.
+The primary perceptual prior knowledge of diffusion models is encapsulated within the U-Net of the diffusion model.
+Customized heads and loss functions offers flexibility and may lead to faster inference speed and improved results.
 3.3 What About the Timesteps and Text Prompts?
@@
 Similarly, the utility of timesteps is reduced, as the single-step paradigm does not require progressive denoising.
-Table 3: Quantitative comparisons among different timesteps and text prompts of diffusion models on depth estimation.
+Table 3: Quantitative comparisons among different timesteps and text prompts on depth estimation.
 Setting Text Prompt Train / Infer KITTI NYU ScanNet DIODE ETH3D
@@
 The significance of the denoiser cannot be overstated.
-However, the strategies for its utilization are worth careful consideration.
+However, the strategies for its utilization are worth a careful study.
 Should we freeze the denoiser, utilize its intermediate features, and merely fine-tune the decoder for specific tasks?
-Alternatively, could we employ LoRA (Hu et al., 2022) instead of extensively fine-tuning the entire denoiser?
+Alternatively, can we employ LoRA (Hu et al., 2022) instead of extensively fine-tuning the entire denoiser?
 Unfortunately, the evidence suggests that neither approach is ideal.
 As illustrated in Table 4, freezing the denoiser significantly compromises performance.
-Although incorporating LoRA offers some advantages, it can not fully leverage the potential of denoiser, especially with regular LoRA ranks of 4 and 16.
+Although incorporating LoRA offers some advantages, it may not fully leverage the potential of denoiser, especially with regular LoRA ranks of 4 and 16.
 This limitation likely stems from the substantial differences between the noise-to-image denoising process and the image-to-perception prediction task.
@@
 Finding 4.
-Fine-tuning the denoiser is preferable compared to either merely utilizing its intermediate features or training it with LoRA.
+Fine-tuning the denoiser appears to be preferable for achieving better results, compared to either merely utilizing its intermediate features or training a LoRA.
 3.5 Is the Training Data Quality Essential?
@@
 This can be done with either a VAE decoder or a customized lightweight decoder.
-Additionally, employing pixel-specific customized loss functions can further enhance detail and overall performance.
+Additionally, employing pixel-specific customized losses can further enhance detail and overall performance.
 We compare our deterministic single-step perception method with previous multi-step paradigms in Fig.
@@
 4.1 Geometric Estimation
+For geometry evaluation, the ensemble size, inference resolution, valid evaluation depth range (specific for depth estimation), and evaluation average paradigm (average by pixels or average by the number of images) can be different for each method.
+To compare these approaches fairly, we follow the open-source evaluation code of Marigold (Ke et al., 2024) for depth and DSINE (Bae & Davison, 2024a) for surface normal, and evaluate the performance of partial existing SOTA methods with their officially released model weights.
+They are labeled with in the Table.
 […]
 Table 6: Quantitative comparison of affine-invariant depth estimation on five zero-shot datasets.
+Part of the reported results () are evaluated following the evaluation protocol of Marigold by ourselves.
 Method Training KITTI NYU ScanNet DIODE ETH3D
@@
 LeReS (Yin et al., 2021) 354K 0.149 0.784 0.090 0.916 0.091 0.917 0.271 0.766 0.171 0.777
-HDN (Zhang et al., 2022a) 300K 0.115 0.867 0.069 0.948 0.080 0.939 0.246 0.780 0.121 0.833
+HDN (Zhang et al., 2022) 300K 0.115 0.867 0.069 0.948 0.080 0.939 0.246 0.780 0.121 0.833
+GeoWizard (Fu et al., 2024b) 280K 0.097 0.921 0.052 0.966 0.061 0.953 0.297 0.792 0.064 0.961
+DepthFM (Gui et al., 2024) 63K 0.083 0.934 0.065 0.956 - - 0.225 0.800 - -
 Marigold (Ke et al., 2024) 74K 0.099 0.916 0.055 0.964 0.064 0.951 0.308 0.773 0.065 0.960
+DMP Official (Lee et al., 2024) - 0.240 0.622 0.109 0.891 0.146 0.814 0.361 0.706 0.128 0.857
 GeoWizard (Fu et al., 2024b) 280K 0.129 0.851 0.059 0.959 0.066 0.953 0.328 0.753 0.077 0.940
@@
 We evaluate mean $\downarrow$, median error $\downarrow$ (med.), and the percentages of pixels $\uparrow$ with five thresholds.
+Part of the reported results () are evaluated following the evaluation protocol of DSINE by ourselves.
 Method Training NYU v2 ScanNet Sintel
@@
 Metric3D v2 (Hu et al., 2024) 8.8M 13.5 6.7 40.1 53.5 65.9 82.6 87.7 11.8 5.5 46.6 60.7 71.6 85.4 89.7 22.8 14.2 18.4 28.5 41.6 66.7 75.8
+Geowizard (Fu et al., 2024b) 280K 17.0 - - - 56.5 - - 15.4 - - - 61.6 - - - - - - - - -
 DINSE (Bae & Davison, 2024a) 160K 16.4 8.4 32.8 46.3 59.6 77.7 83.5 16.2 8.3 29.8 45.9 61.0 78.7 84.4 34.9 28.1 8.9 14.1 21.5 41.5 52.7
@@
 Furthermore, our disparity model (inverse of the depth) shows much better performance on datasets with outdoor scenes, such as KITTI and DIODE, but less performance on indoor datasets.
-By replacing the VAE decoder with a lightweight DPT head, GenPercept can infer faster without bearing the performance penalty.
+Therefore, we suggest adopting the depth model for indoor scenes and the disparity model for outdoor scenes experimentally.
+By replacing the VAE decoder with a lightweight DPT head (Ranftl et al., 2021), GenPercept can infer faster without bearing the performance penalty.
 For surface normal estimation, the image angular loss brings significant performance improvement thanks to our one-step estimation paradigm.
@@
 3.
-We observe excellent generalization of our models in that they can estimate accurate geometric information and promising details not only on diverse real scenes and synthetic scenes, but also on comics, color drafts, and even some sketches.
+We observe excellent generalization of our models in that they can estimate accurate geometric information and promising details not only on diverse real and synthetic scenes, but also on comics, color drafts, and even sketches.
 Figure 3: Qualitative results of monocular depth and surface normal estimation. The model works surprisingly well on out-of-domain images (sketch and cartoon images).
@@
 IS-Net (Qin et al., 2022) 0.791 0.717 0.074 0.813 0.856 1116 0.827 0.753 0.072 0.830 0.870 2888 0.799 0.726 0.070 0.819 0.858 1016
+MVANet (Yu et al., 2024) 0.904 0.861 0.035 0.909 0.937 878 0.911 0.857 0.041 0.903 0.944 2301 0.916 0.855 0.035 0.905 0.938 790
 Our Genpercept 0.857 0.835 0.04 0.87 0.934 1511 0.848 0.829 0.049 0.854 0.938 3799 0.863 0.839 0.039 0.872 0.936 1364
-Table 9: Quantitative results of semantic segmentation performance on Hypersim and ADE20k datasets.
+Our Genpercept (infer. at 1024px) 0.877 0.859 0.035 0.887 0.941 1262 0.874 0.858 0.041 0.874 0.947 3321 0.875 0.856 0.036 0.885 0.939 1176
+Table 9: Quantitative results of semantic segmentation on Hypersim and ADE20k.
 […]
-(Hypersim) (ADE20K)
-Our GenPercept 46.0 34.1
-Our GenPercept 52.9 38.3
-Table 10: Quantitative comparisons of image matting on the P3M-500-NP test set.
-Method SAD $\downarrow$ MAD $\downarrow$ MSE $\downarrow$ CONN $\downarrow$
-HATT (Qiao et al., 2020) 30.35 0.0176 0.0072 27.42
+Dataset (Hypersim) (ADE20K)
+GenPercept (Train UperNet) Hypersim 46.0 34.1
+GenPercept (Train U-Net + UperNet) 52.9 38.3
+Mask2Former R50 ADE20K - 47.2
+Mask2Former Swin-T - 47.7
+Mask2Former Swin-L - 56.4
+GenPercept (Train U-Net + UperNet) - 50.2
+Table 10: Quantitative comparisons of image matting on the P3M-500-NP and AIM500.
+Method Test Dataset SAD $\downarrow$ MAD $\downarrow$ MSE $\downarrow$ CONN $\downarrow$
+HATT (Qiao et al., 2020) P3M-500-NP 30.35 0.0176 0.0072 27.42
 SHM (Chen et al., 2018b) 20.77 0.0122 0.0093 17.09
@@
 P3M-Net (Li et al., 2021) 11.23 0.0065 0.0035 12.51
+ViTAE-S (Ma et al., 2023) 7.59 0.0044 0.0019 6.96
 Our GenPercept 12.77 0.0074 0.0027 10.46
+ViTAE-S (Ma et al., 2023) AIM500 112.52 0.0608 0.0602 43.18
+Our GenPercept (Zero-shot) 75.5 0.0444 0.0242 36.74
 Dichotomous Image Segmentation.
@@
 The pixel-wise MSE loss is utilized during training.
-Quantitative results of dichotomous image segmentation are shown in table 8, respectively.
+Quantitative results of dichotomous image segmentation are shown in Table 8.
 We only show partial results due to paper page limitations, full comparisons are accessible in the supplementary material.
-Our proposed model outperforms all other methods by a significant margin on this challenging dataset across most evaluation metrics.
+GenPercept outperforms methods like HySM (Nirkin et al., 2021) and IS-Net (Qin et al., 2022) on this challenging dataset across most evaluation metrics, but there exists room for further improvement compared to SoTA methods like MVANet (Yu et al., 2024).
 As shown in Fig.
-5, the segmentation results of our approach are more refined, providing a more detailed foreground mask.
+5, our approach provides a detailed foreground mask.
 For thin lines and meticulous objects that are difficult for previous methods to process, our method can also output accurate segmentation results.
@@
 Another choice involves using a customized segmentation head.
-We incorporate a custom segmentation head, namely UperNet (Xiao et al., 2018b), onto the multi-level features extracted by UNet, and the quantitative results are presented in table 10.
-In addition to evaluating on Hypersim (Roberts et al., 2021), we also test the model’s zero-shot performance on a subset of the ADE20k (Zhou et al., 2017) validation set, which contains overlapping classes.
-The results demonstrate the model’s capacity for zero-shot generalization.
+We incorporate a custom segmentation head, namely UperNet (Xiao et al., 2018b), onto the multi-level features extracted by UNet.
+For the UperNet segmentation head, we follow the traditional semantic segmentation format to use n-channel output, where n is the number of categories.
+The quantitative results are presented in Table 10, we test the model’s performance on Hypersim (Roberts et al., 2021) and zero-shot ability on a subset of the ADE20k (Zhou et al., 2017) validation set, which contains overlapping classes.
+Besides, we compare with Mask2Former (Cheng et al., 2022) by training on ADE20K.
+GenPercept outperforms ResNet50 (He et al., 2016) and Swin-T (Liu et al., 2021) of Mask2Former but achieves lower performance than Swin-L (Liu et al., 2021).
 Figure 4: Quanlitative results of semantic segmentation in the wild.
@@
 Automatic image matting seeks to remove this dependency by directly estimating these components from the image alone.
-Implementation Details.
-We utilize P3M10K (Li et al., 2021), the largest portrait matting dataset with high-resolution portrait images along with high-quality alpha masks to train our model.
-The training set contains 9,421 high-quality images and annotations and the test set P3M-500-NP is composed of 500 public images from the Internet.
-We train our GenPercept with pixel-wise MSE loss to further improve the final performance.
+The implementation details can be found in the supplementary material.
 Quantitative and Qualitative Results.
@@
 MSE and CONN focus on L2 distance and connectivity that better reflects human intuition.
-As shown in table 10, our GenPercept is comparable with the state-of-the-art methods.
+As shown in Table 10, our GenPercept is less accurate compared with the state-of-the-art methods.
+However, when transferring the human image matting ability to general image matting tasks, GenPercept achieves much better performance.
+It proves the robustness brought by the prior knowledge of diffusion models pre-trained on the LAION dataset.
 Quantitative results of image matting are shown in Fig.
 6.
+Please see supplementary for more visualization.
 Figure 6: Visualization of image matting on the P3M-500-NP test set.
@@
 By means of contrastive learning, MoCo (He et al., 2020) and CLIP (Radford et al., 2021) acquire rich visual and semantic representations, leveraging their advantages in joint visual and semantic modeling to enhance the performance of multimodal tasks.
-DINO (Zhang et al., 2022b), through self-distillation, endows Vision Transformer and convolutional networks with comparable visual representation quality and demonstrates that self-supervised ViT representations contain explicit semantic segmentation information.
+DINO (Caron et al., 2021), through self-distillation, endows Vision Transformer and convolutional networks with comparable visual representation quality and demonstrates that self-supervised ViT representations contain explicit semantic segmentation information.
+DINOv2 (Oquab et al., 2024) leverages self-supervised learning on a large curated dataset and exhibits remarkable zero-shot generalization capabilities across computer vision tasks at both image and pixel levels, including classification, semantic segmentation, and depth estimation.
 In our work, we leverage Stable Diffusion (Rombach et al., 2022) as a prior for scene understanding and transfer it to various perception tasks.
@@
 In this work, we introduce GenPercept, an embarrassingly straightforward yet powerful approach to re-use the off-the-shelf UNet trained using diffusion processes.
-In essence, we initialize our image perception networks using pre-trained diffusion models trained on the extensive LAION dataset, without making many modifications.
 GenPercept demonstrates the capability to effectively leverage pre-trained diffusion models across a range of downstream dense perception tasks.
 We contend that our proposed methodology provides an efficient and potent paradigm for harnessing the capabilities of pre-trained diffusion models in dense visual perception tasks.
-Supplementary Materials of
-What Matters When Repurposing Diffusion Models for General Dense Perception Tasks?
+For future research, we suggest investigating the impact of scaling up the volume of fine-tuning data and exploring the key components of pre-training by applying alternative self-supervised pre-training methods on the LAION dataset, such as Masked Autoencoders (MAE) or Contrastive Language-Image Pretraining (CLIP).
+It will be helpful to clarify whether the highly detailed visual predictions produced by existing diffusion models are primarily driven by the extensive LAION dataset or the diffusion pretraining paradigm itself.
+Acknowledgments
+Y. Ge is with The University of Adelaide and his contribution was made when visiting Zhejiang University. C. Shen is the corresponding author.
 Appendix A The formulation of Three Different Pipelines
@@
 $\begin{gathered}\mathbf{z}^{(\mathbf{y})}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{z}^{(\mathbf{y})}+\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon},~{}~{}t=[1,...,T],\\ \mathbf{z}_{t}=\text{concat}(\mathbf{z}^{(\mathbf{y})}_{t},\mathbf{z}^{(\mathbf{x})}),\end{gathered}$ (4)
-where $\bar{\alpha}_{t}=\prod_{s=1}^{t}{(1-\beta_{s})}$, and $\beta_{s}$ is sampled from a variance schedule $\{\beta_{t}\in(0,1)\}_{t=1}^{T}$.
-Then, the denoiser $\bm{v}_{\theta}(\cdot,\cdot)$ is enforced to learn the velocity of “v-prediction” Salimans & Ho (2021) from a timestep $t$ and the corresponding input $\mathbf{z}_{t}$.
+where $\bar{\alpha}_{t}=\prod_{s=1}^{t}{(1-\beta_{s})}$, and $\beta_{s}$ is sampled from a variance scheduler $\{\beta_{t}\in(0,1)\}_{t=1}^{T}$.
+The scheduler is parameters with two hyperparameters $\beta_{start}$ and $\beta_{end}$, which defines the $\beta_{t}$ values of t=0 and t=1000, respectively.
+For a casual timestep $s$, $\beta_{s}$ is computed by linearly interpolating between $\sqrt{\beta_{start}}$ and $\sqrt{\beta_{end}}$, then squaring each interpolated value.
+The formulation is as follows.
+$\begin{gathered}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\beta_{s}=(\sqrt{\beta_{start}}+\frac{s}{T}(\sqrt{\beta_{end}}-\sqrt{\beta_{start}}))^{2},~{}~{}s=[1,...,T]}\end{gathered}$ (5)
+Then, the denoiser $\bm{v}_{\theta}(\cdot,\cdot)$ is enforced to learn the “v-prediction” (Salimans & Ho, 2021) from a timestep $t$ and the corresponding input $\mathbf{z}_{t}$.
 During training, the parameters of VAE are frozen, and only the denoiser $\bm{v}_{\theta}$ is fine-tuned.
@@
 After repeating it for $T$ times, the predicted clean latent code $\hat{\mathbf{z}}^{(\mathbf{y})}_{1\rightarrow 0}$ is computed and sent to the VAE decoder $\mathcal{D}$ and estimate the target label $\hat{\mathbf{y}}$.
-During inference, $m$ randomly sampled noises are introduced to estimate $m$ different predictions, and they are averaged with an ensemble process to reduce the randomness of prediction and improve performance.
+During inference, $m$ randomly sampled noises are introduced to estimate $m$ different predictions, and they are averaged with an ensemble process (Ke et al., 2024) to reduce the randomness of prediction and improve performance.
 A.2 Deterministic Multi-Step Generation
@@
 $\begin{gathered}\mathbf{z}_{t}=\mathbf{z}^{(\mathbf{y})}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{z}^{(\mathbf{y})}+\sqrt{1-\bar{\alpha}_{t}}\mathbf{z}^{(\mathbf{x})},~{}~{}t=[1,...,T],\\ \end{gathered}$ (8)
-Figure 7: Qualitative visualization of ablation study.
-Figure 8: Visualization of different noise forms and proportions in the forward diffusion process.
+Table 11: Runtime comparison of three diffusion for perception pipelines on an RTX 4090 GPU.
+Experimental Setting Ensemble Denoise Steps Inference Time GPU Memory
+Stochastic Multi-step Generation (w. ensemble) 10 10 $\sim$5.74s 16GB
+Stochastic Multi-step Generation (w/o ensemble) 1 10 $\sim$0.79s 6.95GB
+Deterministic Multi-step Generation 1 10 $\sim$0.79s 6.95GB
+Deterministic One-step Inference (Ours) 1 1 $\sim$0.34s 6.95GB
+Deterministic One-step Inference + DPT head (Ours) 1 1 $\sim$0.24s 6.32GB
+Metric3Dv2 (Hu et al., 2024) 1 1 $\sim$0.25s 2.63GB
+DepthAnythingv2 (Yang et al., 2024b) 1 1 $\sim$0.07s 2.82GB
+DSINE (Bae & Davison, 2024a) 1 1 $\sim$0.18s 2.23GB
+Marigold (Ke et al., 2024) 1 10 $\sim$0.79s 6.95GB
+GeoWizard (Fu et al., 2024b) 1 1 $\sim$1.32s 6.81GB
+DepthFM (Gui et al., 2024) 1 2 $\sim$0.41s 6.97GB
+Our GenPercept (DPT head) 1 1 $\sim$0.24s 6.32GB
+Table 12: The impact of training data volume on affine-invariant monocular depth estimation.
+Amount of Data KITTI NYU ScanNet DIODE ETH3D
+[…]
+90K (1/1) 0.100 0.902 0.053 0.966 0.059 0.961 0.309 0.768 0.068 0.956
+45K (1/2) 0.101 0.902 0.056 0.964 0.058 0.963 0.311 0.764 0.070 0.955
+22.5K (1/4) 0.109 0.884 0.056 0.963 0.059 0.963 0.322 0.754 0.073 0.950
+11.2K (1/8) 0.117 0.866 0.060 0.962 0.065 0.957 0.324 0.753 0.076 0.943
+5.6K (1/16) 0.117 0.868 0.063 0.958 0.068 0.952 0.331 0.744 0.084 0.932
 Furthermore, the input latent has been modified to the latent code of input image $\mathbf{z}^{(\mathbf{x})}$ instead of random Gaussian.
@@
 A.3 GenPercept: Deterministic One-step Perception
-In the main paper, we suppose that blending with the RGB image is much simpler compared to Gaussian noise, and we increase the blending proportion of the RGB image to make it more difficult.
-This is realized by increasing the initial value $\beta_{0}$ and final values $\beta_{T}$ of the DDPM schedule.
-When all the beta values $\{\beta_{t}\}_{t=1}^{T}$ are increased to 1, and $\bar{\alpha}_{t}=\prod_{s=1}^{t}{(1-\beta_{s})}=0$, the formulation of eq.
+In the main paper, We set the ($\beta_{start}$, $\beta_{end}$) values to 1, and $\bar{\alpha}_{t}=\prod_{s=1}^{t}{(1-\beta_{s})}=0$, the formulation of Eq.
 8 to Eq.
@@
 $\begin{gathered}\mathcal{L}=\mathbb{E}_{(\mathbf{z}^{(\mathbf{x})},\mathbf{z}^{(\mathbf{y})})}\left\|-\mathbf{z}^{(\mathbf{y})}-\bm{v}_{\theta}(\mathbf{z}^{(\mathbf{x})},t=1)\right\|^{2}_{2},\\ \hat{\mathbf{z}}^{(\mathbf{y})}_{1\rightarrow 0}=-\bm{v}_{\theta}(\mathbf{z}^{(\mathbf{x})},t=1),~{}~{}\hat{\mathbf{y}}=\mathcal{D}(\hat{\mathbf{z}}^{(\mathbf{y})}_{1\rightarrow 0}).\end{gathered}$ (12)
-Appendix B More Visualization Analysis
+Appendix B A Difference Between Image Generation and Visual Perception
+In text-guided image generation, a single textual input can correspond to an immense variety of potential images.
+This inherent uncertainty makes generating a high-quality image directly from random noise in a single step extremely challenging.
+Therefore, the multi-step generation enables the model to incrementally remove noise, progressively refining details and textures at each stage, which effectively simplifies the task.
+However, visual perception tasks conditioned on an RGB image are deterministic without any randomness, and such an easy injective mapping can be estimated with a one-step inference process, as most of the traditional visual perception methods do.
+While Marigold series algorithms aim to leverage diffusion models’ ability of generating highly detailed images to enhance visual perception with precise details, reformulating straightforward deterministic tasks as a denoising process can further simplify this problem, enforcing the network to exploit ”shortcuts”, as described in Section 3.1 of the main paper and illustrated in Fig.
+8.
+Appendix C Runtime Analysis
+In this section, we quantitatively analyze the inference times of the three aforementioned pipelines, as summarized in Table 11.
+The runtime is evaluated by averaging the inference times over 100 images with a resolution of $768\times 768$, using an RTX 4090 GPU.
+For ”Stochastic Multi-step Generation” methods (Ke et al., 2024; Fu et al., 2024b; Gui et al., 2024), such as Marigold (Ke et al., 2024), they rely on an ensemble process where multiple inferences are performed with varying random noise inputs to mitigate uncertainties introduced by Gaussian noise.
+Consequently, this approach is computationally expensive.
+On the other hand, ”Deterministic Multi-step Generation” methods (Lee et al., 2024) involve multiple denoising steps, which significantly reduce inference efficiency.
+In contrast, our proposed one-step inference paradigm demonstrates a runtime that is 94% and 57% less than those of multi-step methods with ensemble and without ensemble, respectively.
+Furthermore, by incorporating a customized head, such as the DPT head (Ranftl et al., 2021), both runtime and GPU memory requirements are further reduced by 27% without compromising performance, maintaining a competitive level of accuracy and robustness.
+Compared to existing state-of-the-art diffusion-based methods, our proposed GenPercept achieves a notable improvement in inference speed, attributed to the innovative one-step inference paradigm and the customized head.
+While our method demonstrates inference speeds comparable to Metric3Dv2 (Hu et al., 2024) and DSINE (Bae & Davison, 2024a), it falls behind DepthAnythingV2 (Yang et al., 2024b).
+Note that the superior performance of DepthAnythingV2 is facilitated by its training on a relatively lightweight model, bolstered by extensive labeled and unlabeled datasets, and supported by substantial computational resources distributed across multiple GPUs.
+Figure 7: Qualitative visualization of ablation study.
+Figure 8: Visualization of different noise forms and proportions in the forward diffusion process.
+Appendix D The Impact of Data Volume
+With only around 50K Hypersim (Roberts et al., 2021) and 40K Virtual KITTI (Cabon et al., 2020) fine-tuning data, GenPercept can generalize well to diverse tasks and datasets.
+How much data is needed for transferring at least?
+We gradually reduce the amount of training data.
+As shown in Table 12, less training data results in slightly worse performance, but GenPercept still shows much robustness to the data volume.
+Appendix E More Visualization Analysis
 Quantitative Comparisons of Ablations.
@@
 7.
-The estimated depth detail remains comparable with a customized “DPT head”.
+The estimated depth detail remains comparable with a customized “DPT head” (Ranftl et al., 2021).
 Models trained with lower-quality data like “Taskonomy + Cityscapes” or without the pre-trained VAE decoder parameters suffer from a decline in the ability to predict details.
@@
 8.
-With larger $\beta$ values of the DDIM scheduler, the proportion of noise will be much higher during the forward diffusion process.
-When $\beta_{start},\beta_{end}$ reaches (1.0, 1.0), the noisy latent will be pure noise latent, which is Gaussian noise and RGB latent for (a) and (b), respectively.
-Appendix C An Extra Attempt on Human Pose Estimation
+With larger ($\beta_{start}$, $\beta_{end}$) values of the DDIM scheduler, the proportion of noise will be much higher during the forward diffusion process.
+When ($\beta_{start}$, $\beta_{end}$) reaches (1.0, 1.0), the noisy latent will be pure noise latent, which is Gaussian noise and RGB latent for (a) and (b), respectively.
+Quantitative Comparisons of Generalization.
+We compare the generalization performance of models trained on synthetic (50K Hypersim (Roberts et al., 2021) + 40K Virtual KITTI (Cabon et al., 2020)) and real data (50K Taskonomy (Zamir et al., 2018) + 40K Cityscapes (Cordts et al., 2016)) for out-of-distribution scenarios.
+As illustrated in Fig.
+9, the robustness of GenPercept trained on real data is comparable to that trained on synthetic data in challenging scenes, such as underwater environments, non-realistic renderings, and evening settings.
+Notably, models trained on synthetic data demonstrate superior accuracy in estimating transparent objects and capturing geometric details, owing to the high-quality and densely labeled synthetic ground truth.
+Figure 9: Quantitative comparisons of generalization for models trained on synthetic and real data.
+Appendix F An Extra Attempt on Human Pose Estimation
 Task Definition Human Pose Estimation is a task aimed at determining the spatial configuration of a person or object in a given image or video.
@@
 As shown in Fig.
-9, it is generalizable to unseen objects in the training set.
+10, it is generalizable to unseen objects in the training set.
 To evaluate the performance on COCO, we use the customized keypoint head of ViTPose (Xu et al., 2022) for decoding the output.
@@
 Figure 10: Generalized test results of keypoint detection.
-Appendix D Implementation Details
-Implementation Details of Preliminary Section in the Main Paper.
-Unless specified otherwise, we train for 30000 iterations to estimate the affine-invariant monocular depth on a mixture of the Hypersim (Roberts et al., 2021) (90%) and Virtual KITTI (Cabon et al., 2020) (10%) datasets.
+Appendix G Implementation Details
+Implementation Details of Exploration Experiments in Section 3 of the Main Paper.
+Unless specified otherwise, we follow Marigold’s training setting and train for 30000 iterations to estimate the affine-invariant monocular depth.
+The training dataset contains 50K Hypersim (Roberts et al., 2021) images and 40K Virtual KITTI (Cabon et al., 2020) images, and these images are sampled with a sampling rate of 90% for Hypersim and 10% for Virtual KITTI.
 We freeze the VAE AutoEncoder and fine-tune the U-Net of Stable Diffusion v2.1 to estimate the ground-truth label latent, with a resolution of (480, 640), a batch size of 32, and a learning rate of 3e-5.
@@
 Evaluation results of absolute relative error (AbsRel) and $\delta_{1}$ are reported on five unseen monocular depth datasets, including KITTI (Geiger et al., 2013), NYU (Silberman et al., 2012), ScanNet (Dai et al., 2017), DIODE (Vasiljevic et al., 2019), and ETH3D (Schops et al., 2017).
-Appendix E More Quantitative and Qualitive Experiments
-E.1 Monocular Depth Estimation
+Appendix H More Quantitative and Qualitive Experiments
+H.1 Monocular Depth Estimation
 More Qualitative Evaluation.
@@
 Compared to Marigold (Ke et al., 2024), our method achieves better relative depth visualization.
-E.2 Surface Normal Estimation
+H.2 Surface Normal Estimation
 More Qualitative Evaluation.
@@
 Our GenPercept is only trained on one synthetic dataset (Hypersim (Roberts et al., 2021)), and can estimate much more detailed surface normal maps.
-E.3 Dichotomous Image Segmentation
+H.3 Dichotomous Image Segmentation
 More Quantitative Evaluation.
-To conduct a comprehensive evaluation, we compare our approach with numerous previous models including models for medical image segmentation (Ronneberger et al., 2015), salient object detection (Qin et al., 2019; Zhao et al., 2020; Wei et al., 2020; Chen et al., 2020; Qin et al., 2020), camouflaged object detection (Fan et al., 2021a; Mei et al., 2021), semantic segmentation (Zhao et al., 2017; Chen et al., 2018a; Wang et al., 2020; Yu et al., 2018; Zhao et al., 2018; Howard et al., 2019; Fan et al., 2021c; Nirkin et al., 2021) and the model IS-Net (Qin et al., 2022) specifically trained for DIS.
-As shown in table 12, our proposed model significantly outperforms other methods across most evaluation metrics on this challenging dataset.
-The results highlight the effectiveness of our approach for DIS.
+To conduct a comprehensive evaluation, we compare our approach with numerous previous models including models for medical image segmentation (Ronneberger et al., 2015), salient object detection (Qin et al., 2019; Zhao et al., 2020; Wei et al., 2020; Chen et al., 2020; Qin et al., 2020), camouflaged object detection (Fan et al., 2021a; Mei et al., 2021), semantic segmentation (Zhao et al., 2017; Chen et al., 2018a; Wang et al., 2020; Yu et al., 2018; Zhao et al., 2018; Howard et al., 2019; Fan et al., 2021c; Nirkin et al., 2021) and models like IS-Net (Qin et al., 2022) and MVANet (Yu et al., 2024) specifically trained for DIS.
+As shown in Table 14, our proposed model significantly outperforms methods like IS-Net across most evaluation metrics on this challenging dataset.
+Compared with MVANet, which inferences with a resolution of $1024\times 1024$, our GenPercept achieves slightly lower performance, but the results remains competitive and can highlight the effectiveness of our approach for DIS.
 Table 14: Quantitative results on DIS5K validation and testing sets.
@@
 IS-Net
+MVANet
 Ours
-DIS-VD $maxF_{\beta}\uparrow$ 0.692 0.731 0.678 0.685 0.648 0.748 0.665 0.691 0.691 0.660 0.726 0.662 0.697 0.714 0.696 0.734 0.791 0.857
-$F^{w}_{\beta}\uparrow$ 0.586 0.641 0.574 0.595 0.542 0.656 0.584 0.604 0.603 0.568 0.641 0.548 0.609 0.642 0.613 0.640 0.717 0.835
-${}M~{}\downarrow$ 0.113 0.094 0.110 0.107 0.118 0.090 0.110 0.106 0.102 0.114 0.095 0.116 0.102 0.092 0.103 0.096 0.074 0.040
-$S_{\alpha}\uparrow$ 0.745 0.768 0.723 0.733 0.718 0.781 0.727 0.740 0.744 0.716 0.767 0.728 0.747 0.758 0.740 0.773 0.813 0.870
-$E_{\phi}^{m}\uparrow$ 0.785 0.816 0.783 0.800 0.765 0.823 0.798 0.811 0.802 0.796 0.824 0.767 0.811 0.841 0.817 0.814 0.856 0.934
-${\rm HCE}_{\gamma}\downarrow$ 1337 1402 1493 1567 1555 1413 1568 1606 1588 1520 1560 1660 1503 1625 1598 1324 1116 1511
-DIS-TE1 $maxF_{\beta}\uparrow$ 0.625 0.688 0.620 0.640 0.598 0.694 0.644 0.646 0.645 0.601 0.668 0.595 0.631 0.669 0.648 0.695 0.740 0.841
-$F^{w}_{\beta}\uparrow$ 0.514 0.595 0.517 0.549 0.495 0.601 0.558 0.552 0.557 0.506 0.579 0.474 0.535 0.595 0.562 0.597 0.662 0.814
-${}M~{}\downarrow$ 0.106 0.084 0.099 0.095 0.103 0.083 0.094 0.094 0.089 0.102 0.088 0.108 0.095 0.083 0.090 0.082 0.074 0.038
-$S_{\alpha}\uparrow$ 0.716 0.754 0.701 0.721 0.705 0.760 0.727 0.722 0.725 0.694 0.742 0.695 0.716 0.740 0.723 0.761 0.787 0.868
-$E_{\phi}^{m}\uparrow$ 0.750 0.801 0.766 0.783 0.750 0.801 0.791 0.786 0.791 0.772 0.797 0.741 0.784 0.818 0.798 0.803 0.820 0.918
-${\rm HCE}_{\gamma}\downarrow$ 233 220 230 244 271 224 274 253 267 234 262 288 234 274 249 205 149 204
-DIS-TE2 $maxF_{\beta}\uparrow$ 0.703 0.755 0.702 0.712 0.673 0.756 0.700 0.720 0.724 0.681 0.747 0.680 0.716 0.743 0.720 0.759 0.799 0.876
-$F^{w}_{\beta}\uparrow$ 0.597 0.668 0.598 0.620 0.570 0.668 0.618 0.633 0.636 0.587 0.664 0.564 0.627 0.672 0.636 0.667 0.728 0.852
-${}M~{}\downarrow$ 0.107 0.084 0.102 0.097 0.109 0.085 0.099 0.096 0.092 0.105 0.087 0.111 0.095 0.083 0.092 0.085 0.070 0.035
-$S_{\alpha}\uparrow$ 0.755 0.786 0.737 0.755 0.735 0.788 0.753 0.761 0.763 0.729 0.784 0.740 0.759 0.777 0.759 0.794 0.823 0.884
-$E_{\phi}^{m}\uparrow$ 0.796 0.836 0.804 0.820 0.786 0.833 0.823 0.829 0.828 0.813 0.840 0.781 0.826 0.856 0.834 0.832 0.858 0.938
-${\rm HCE}_{\gamma}\downarrow$ 474 480 501 542 574 490 593 567 586 516 555 621 512 600 556 451 340 480
-DIS-TE3 $maxF_{\beta}\uparrow$ 0.748 0.785 0.726 0.743 0.699 0.798 0.730 0.751 0.747 0.717 0.784 0.710 0.752 0.772 0.745 0.792 0.830 0.885
-$F^{w}_{\beta}\uparrow$ 0.644 0.696 0.620 0.656 0.590 0.707 0.641 0.664 0.657 0.623 0.700 0.595 0.664 0.702 0.662 0.701 0.758 0.862
-${}M~{}\downarrow$ 0.098 0.083 0.103 0.092 0.109 0.079 0.096 0.092 0.092 0.102 0.080 0.109 0.091 0.078 0.090 0.079 0.064 0.035
-$S_{\alpha}\uparrow$ 0.780 0.798 0.747 0.773 0.748 0.809 0.766 0.777 0.774 0.749 0.805 0.757 0.780 0.794 0.771 0.811 0.836 0.883
-$E_{\phi}^{m}\uparrow$ 0.827 0.856 0.815 0.848 0.801 0.858 0.849 0.854 0.843 0.833 0.869 0.801 0.852 0.880 0.855 0.857 0.883 0.951
-${\rm HCE}_{\gamma}\downarrow$ 883 948 972 1059 1058 965 1096 1082 $1$ 999 1049 1146 1001 1136 1081 887 687 973
-DIS-TE4 $maxF_{\beta}\uparrow$ 0.759 0.780 0.729 0.721 0.670 0.795 0.699 0.731 0.725 0.715 0.772 0.710 0.749 0.736 0.731 0.782 0.827 0.848
-$F^{w}_{\beta}\uparrow$ 0.659 0.693 0.625 0.633 0.559 0.705 0.616 0.647 0.630 0.621 0.687 0.598 0.663 0.664 0.652 0.693 0.753 0.829
-${}M~{}\downarrow$ 0.102 0.091 0.109 0.107 0.127 0.087 0.113 0.107 0.107 0.111 0.092 0.114 0.099 0.098 0.102 0.091 0.072 0.049
-$S_{\alpha}\uparrow$ 0.784 0.794 0.743 0.752 0.723 0.807 0.744 0.763 0.758 0.744 0.792 0.755 0.776 0.770 0.762 0.802 0.830 0.854
-$E_{\phi}^{m}\uparrow$ 0.821 0.848 0.803 0.825 0.767 0.847 0.824 0.838 0.815 0.820 0.854 0.788 0.837 0.848 0.841 0.842 0.870 0.938
-${\rm HCE}_{\gamma}\downarrow$ 3218 3601 3654 3760 3678 3653 3683 3803 3806 3709 3864 3999 3690 3817 3819 3331 2888 3799
-Overall DIS-TE (1-4) $maxF_{\beta}\uparrow$ 0.708 0.752 0.694 0.704 0.660 0.761 0.693 0.712 0.710 0.678 0.743 0.674 0.711 0.729 0.710 0.757 0.799 0.863
-$F^{w}_{\beta}\uparrow$ 0.603 0.663 0.590 0.614 0.554 0.670 0.608 0.624 0.620 0.584 0.658 0.558 0.622 0.658 0.628 0.665 0.726 0.839
-${}M~{}\downarrow$ 0.103 0.086 0.103 0.098 0.112 0.083 0.101 0.097 0.095 0.105 0.087 0.110 0.095 0.085 0.094 0.084 0.070 0.039
-$S_{\alpha}\uparrow$ 0.759 0.783 0.732 0.750 0.728 0.791 0.747 0.756 0.755 0.729 0.781 0.737 0.758 0.770 0.754 0.792 0.819 0.872
-$E_{\phi}^{m}\uparrow$ 0.798 0.835 0.797 0.819 0.776 0.835 0.822 0.827 0.819 0.810 0.840 0.778 0.825 0.850 0.832 0.834 0.858 0.936
-${\rm HCE}_{\gamma}\downarrow$ 1202 1313 1339 1401 1395 1333 1411 1427 1442 1365 1432 1513 1359 1457 1426 1218 1016 1364
-More Qualitative Evaluation.
+Ours (1024px)
+DIS-VD $maxF_{\beta}\uparrow$ 0.692 0.731 0.678 0.685 0.648 0.748 0.665 0.691 0.691 0.660 0.726 0.662 0.697 0.714 0.696 0.734 0.791 0.904 0.857 0.877
+$F^{w}_{\beta}\uparrow$ 0.586 0.641 0.574 0.595 0.542 0.656 0.584 0.604 0.603 0.568 0.641 0.548 0.609 0.642 0.613 0.640 0.717 0.861 0.835 0.859
+${}M~{}\downarrow$ 0.113 0.094 0.110 0.107 0.118 0.090 0.110 0.106 0.102 0.114 0.095 0.116 0.102 0.092 0.103 0.096 0.074 0.035 0.040 0.035
+$S_{\alpha}\uparrow$ 0.745 0.768 0.723 0.733 0.718 0.781 0.727 0.740 0.744 0.716 0.767 0.728 0.747 0.758 0.740 0.773 0.813 0.909 0.870 0.887
+$E_{\phi}^{m}\uparrow$ 0.785 0.816 0.783 0.800 0.765 0.823 0.798 0.811 0.802 0.796 0.824 0.767 0.811 0.841 0.817 0.814 0.856 0.937 0.934 0.941
+${\rm HCE}_{\gamma}\downarrow$ 1337 1402 1493 1567 1555 1413 1568 1606 1588 1520 1560 1660 1503 1625 1598 1324 1116 878 1511 1262
+DIS-TE1 $maxF_{\beta}\uparrow$ 0.625 0.688 0.620 0.640 0.598 0.694 0.644 0.646 0.645 0.601 0.668 0.595 0.631 0.669 0.648 0.695 0.740 0.893 0.841 0.850
+$F^{w}_{\beta}\uparrow$ 0.514 0.595 0.517 0.549 0.495 0.601 0.558 0.552 0.557 0.506 0.579 0.474 0.535 0.595 0.562 0.597 0.662 0.823 0.814 0.827
+${}M~{}\downarrow$ 0.106 0.084 0.099 0.095 0.103 0.083 0.094 0.094 0.089 0.102 0.088 0.108 0.095 0.083 0.090 0.082 0.074 0.037 0.038 0.036
+$S_{\alpha}\uparrow$ 0.716 0.754 0.701 0.721 0.705 0.760 0.727 0.722 0.725 0.694 0.742 0.695 0.716 0.740 0.723 0.761 0.787 0.879 0.868 0.878
+$E_{\phi}^{m}\uparrow$ 0.750 0.801 0.766 0.783 0.750 0.801 0.791 0.786 0.791 0.772 0.797 0.741 0.784 0.818 0.798 0.803 0.820 0.911 0.918 0.919
+${\rm HCE}_{\gamma}\downarrow$ 233 220 230 244 271 224 274 253 267 234 262 288 234 274 249 205 149 103 204 165
+DIS-TE2 $maxF_{\beta}\uparrow$ 0.703 0.755 0.702 0.712 0.673 0.756 0.700 0.720 0.724 0.681 0.747 0.680 0.716 0.743 0.720 0.759 0.799 0.925 0.876 0.880
+$F^{w}_{\beta}\uparrow$ 0.597 0.668 0.598 0.620 0.570 0.668 0.618 0.633 0.636 0.587 0.664 0.564 0.627 0.672 0.636 0.667 0.728 0.874 0.852 0.859
+${}M~{}\downarrow$ 0.107 0.084 0.102 0.097 0.109 0.085 0.099 0.096 0.092 0.105 0.087 0.111 0.095 0.083 0.092 0.085 0.070 0.03 0.035 0.034
+$S_{\alpha}\uparrow$ 0.755 0.786 0.737 0.755 0.735 0.788 0.753 0.761 0.763 0.729 0.784 0.740 0.759 0.777 0.759 0.794 0.823 0.915 0.884 0.892
+$E_{\phi}^{m}\uparrow$ 0.796 0.836 0.804 0.820 0.786 0.833 0.823 0.829 0.828 0.813 0.840 0.781 0.826 0.856 0.834 0.832 0.858 0.944 0.938 0.938
+${\rm HCE}_{\gamma}\downarrow$ 474 480 501 542 574 490 593 567 586 516 555 621 512 600 556 451 340 246 480 410
+DIS-TE3 $maxF_{\beta}\uparrow$ 0.748 0.785 0.726 0.743 0.699 0.798 0.730 0.751 0.747 0.717 0.784 0.710 0.752 0.772 0.745 0.792 0.830 0.936 0.885 0.898
+$F^{w}_{\beta}\uparrow$ 0.644 0.696 0.620 0.656 0.590 0.707 0.641 0.664 0.657 0.623 0.700 0.595 0.664 0.702 0.662 0.701 0.758 0.89 0.862 0.879
+${}M~{}\downarrow$ 0.098 0.083 0.103 0.092 0.109 0.079 0.096 0.092 0.092 0.102 0.080 0.109 0.091 0.078 0.090 0.079 0.064 0.031 0.035 0.032
+$S_{\alpha}\uparrow$ 0.780 0.798 0.747 0.773 0.748 0.809 0.766 0.777 0.774 0.749 0.805 0.757 0.780 0.794 0.771 0.811 0.836 0.92 0.883 0.896
+$E_{\phi}^{m}\uparrow$ 0.827 0.856 0.815 0.848 0.801 0.858 0.849 0.854 0.843 0.833 0.869 0.801 0.852 0.880 0.855 0.857 0.883 0.954 0.951 0.954
+${\rm HCE}_{\gamma}\downarrow$ 883 948 972 1059 1058 965 1096 1082 $1$ 999 1049 1146 1001 1136 1081 887 687 512 973 809
+DIS-TE4 $maxF_{\beta}\uparrow$ 0.759 0.780 0.729 0.721 0.670 0.795 0.699 0.731 0.725 0.715 0.772 0.710 0.749 0.736 0.731 0.782 0.827 0.911 0.848 0.874
+$F^{w}_{\beta}\uparrow$ 0.659 0.693 0.625 0.633 0.559 0.705 0.616 0.647 0.630 0.621 0.687 0.598 0.663 0.664 0.652 0.693 0.753 0.857 0.829 0.858
+${}M~{}\downarrow$ 0.102 0.091 0.109 0.107 0.127 0.087 0.113 0.107 0.107 0.111 0.092 0.114 0.099 0.098 0.102 0.091 0.072 0.041 0.049 0.041
+$S_{\alpha}\uparrow$ 0.784 0.794 0.743 0.752 0.723 0.807 0.744 0.763 0.758 0.744 0.792 0.755 0.776 0.770 0.762 0.802 0.830 0.903 0.854 0.874
+$E_{\phi}^{m}\uparrow$ 0.821 0.848 0.803 0.825 0.767 0.847 0.824 0.838 0.815 0.820 0.854 0.788 0.837 0.848 0.841 0.842 0.870 0.944 0.938 0.947
+${\rm HCE}_{\gamma}\downarrow$ 3218 3601 3654 3760 3678 3653 3683 3803 3806 3709 3864 3999 3690 3817 3819 3331 2888 2301 3799 3321
+Overall DIS-TE (1-4) $maxF_{\beta}\uparrow$ 0.708 0.752 0.694 0.704 0.660 0.761 0.693 0.712 0.710 0.678 0.743 0.674 0.711 0.729 0.710 0.757 0.799 0.916 0.863 0.875
+$F^{w}_{\beta}\uparrow$ 0.603 0.663 0.590 0.614 0.554 0.670 0.608 0.624 0.620 0.584 0.658 0.558 0.622 0.658 0.628 0.665 0.726 0.855 0.839 0.856
+${}M~{}\downarrow$ 0.103 0.086 0.103 0.098 0.112 0.083 0.101 0.097 0.095 0.105 0.087 0.110 0.095 0.085 0.094 0.084 0.070 0.035 0.039 0.036
+$S_{\alpha}\uparrow$ 0.759 0.783 0.732 0.750 0.728 0.791 0.747 0.756 0.755 0.729 0.781 0.737 0.758 0.770 0.754 0.792 0.819 0.905 0.872 0.885
+$E_{\phi}^{m}\uparrow$ 0.798 0.835 0.797 0.819 0.776 0.835 0.822 0.827 0.819 0.810 0.840 0.778 0.825 0.850 0.832 0.834 0.858 0.938 0.936 0.939
+${\rm HCE}_{\gamma}\downarrow$ 1202 1313 1339 1401 1395 1333 1411 1427 1442 1365 1432 1513 1359 1457 1426 1218 1016 790 1364 1176
+More Qualitative Evaluation. As shown in Fig. 14, our method yields refined segmentation results, providing cleaner foreground masks. It also can produce precise outputs for intricate lines.
+Cross Dataset Evaluation.
+To test the generalization ability of our model, we randomly select some images from other datasets(Agustsson & Timofte, 2017; Lin et al., 2014; Shao et al., 2019) and in-the-wild images for experiments.
 As shown in  Fig.
-13, our method yields more refined segmentation results, providing cleaner foreground masks.
-It also produces precise outputs for intricate lines that are challenging for previous approaches to handle.
-Cross Dataset Evaluation.
-To test the generalization ability of our model, we randomly select some images from other datasets(Agustsson & Timofte, 2017; Lin et al., 2014; Shao et al., 2019) for experiments.
-As shown in Fig.
-10, compared to IS-Net and IS-Net-General-UseQin et al.
-(2022), our approach exhibits finer segmentation quality across diverse images, providing cleaner foreground masks.
+13, compared to IS-Net and IS-Net-General-Use(Qin et al., 2022), our approach exhibits finer segmentation quality across diverse images, providing cleaner foreground masks.
+Compared to MVANet (Yu et al., 2024), GenPercept exhibits enhanced robustness when applied to in-the-wild images.
+This improvement can be attributed to its large-scale pre-training on the LAION dataset and the extensive parameterization of the diffusion model.
+In contrast, the backbone of MVANet is pre-trained on the ImageNet dataset (Deng et al., 2009), which may limit its performance in more diverse environments.
 It is noteworthy that IS-Net-General-Use is fine-tuned on extra datasets to enhance generalization, which indicates that our method has a stronger generalization ability.
-Figure 10: Cross dataset comparison of our model and other models.
-E.4 Image Matting
+H.4 Image Matting
+Implementation Details.
+We utilize P3M10K (Li et al., 2021), the largest portrait matting dataset with high-resolution portrait images along with high-quality alpha masks to train our model.
+The training set contains 9,421 high-quality images and annotations and the test set P3M-500-NP is composed of 500 public images from the Internet.
+We train our GenPercept with pixel-wise MSE loss to further improve the final performance.
 More Qualitative Evaluation.
 In Fig.
-14, we showcase more qualitative results for the image matting task.
+15, we showcase more qualitative results for the image matting task.
 It is worth noting that our model works well in various resolutions, light environments, human poses, and human orientations.
-Please zoom in for better visualization and more details.
-E.5 Image Segmentation
+More importantly, our GenPercept model trained on human matting images shows much more robustness to other objects compared to existing P3M10K (Li et al., 2021) SOTA method ViTAE-S (Ma et al., 2023), as illustrated in Fig.
+16.
+It shows that the ViTAE-S overfits the human matting task, while GenPercept preserves the generalization ability.
+Besides, we also train GenPercept on a more general image matting task on the Composition-1K (Xu et al., 2017) dataset.
+As shown in Fig.
+17, GenPercept shows robustness on more types of objects such as semi-transparent objects, hollow objects, etc.
+H.5 Image Segmentation
 More Qualitative Evaluation.
 In Fig.
-15, we showcase more qualitative results for the image segmentation task.
+18, we showcase more qualitative results for the image segmentation task.
 Our method shows much robustness on the trained categories of complex in-the-wild images.
@@
 Please zoom in for better visualization and more details.
-E.6 Human Pose Estimation
+H.6 Human Pose Estimation
 More Qualitative Evaluation.
-In Fig.16, we showcase more qualitative results for the human pose estimation.
-We conduct experiments on MHP datasetLi et al.
-(2017), and we use mmposeContributors (2020) to render the human pose following the setting of Bai et al.
-(2023).
+In Fig.
+19, we showcase more qualitative results for the human pose estimation.
+We conduct experiments on MHP dataset (Li et al., 2017), and we use mmpose (Contributors, 2020) to render the human pose following the setting of (Bai et al., 2023).
 Figure 11: More qualitative results for monocular depth estimation.
 Figure 12: More qualitative comparisons of surface normal estimation.
-Our GenPercept can achieve more detailed results, even compared to the competitive CVPR2024 method DSINE Bae & Davison (2024b).
+Our GenPercept can achieve more detailed results, even compared to the competitive CVPR2024 method DSINE (Bae & Davison, 2024b).
 Note that these two visualization coordinate systems are slightly different.
+Figure 13: Cross dataset comparison of our model and other models.
 Figure 14: Qualitative results of dichotomous image segmentation.
 Figure 15: More qualitative results for image matting.
-Figure 15: More qualitative results for image matting.
+Figure 16: Generalization ability of the human matting model to general image matting images.
+Figure 17: More types of image matting such as semi-transparent objects.
+Figure 18: More qualitative results for image segmentation.
 Figure 19: More qualitative results for human pose estimation. (Left: original Image, Mid: prediction, Right: ground truth)